### v1

# Graph Reinforcement Learning in Power Grids: A Survey

Abstract
The challenges posed by renewable energy and distributed electricity generation motivate the development of deep learning approaches to overcome the lack of flexibility of traditional methods in power grids use cases. The application of Graph Neural Networks is particularly promising due to their ability to learn from graph-structured data present in power grids. Combined with Reinforcement Learning (RL), they can serve as control approaches to determine remedial grid actions. This review analyses the ability of Graph Reinforcement Learning (GRL) to capture the inherent graph structure of power grids to improve representation learning and decision making in different power grid use cases. It distinguishes between common problems in transmission and distribution grids and explores the synergy between RL and GNNs. In transmission grids, GRL typically addresses automated grid management and topology control, whereas on the distribution side, GRL concentrates more on voltage regulation. We analyzed the selected papers based on their graph structure and GNN model, the applied RL algorithm, and their overall contributions. Although GRL demonstrate adaptability in the face of unpredictable events and noisy or incomplete data, it primarily serves as a proof of concept at this stage. There are multiple open challenges and limitations that need to be addressed when considering the application of RL to real power grid operation.
Keywords Power Grids  $\cdot$ Deep Learning  $\cdot$ Deep Reinforcement Learning  $\cdot$ Graph Neural Networks  $\cdot$ Graph Reinforcement Learning
1 Introduction
The role of electrical grid operators, both for transmission and distribution grids, is to ensure cost-efficient availability at all times. However, power systems worldwide are undergoing a paradigm shift driven by the need for CO2 neutrality. The incorporation of renewable distributed generation, and additional load demand due to heating and traffic sector electrification introduce complexities that traditional power system operation struggles to address. These trends necessitate advanced methodologies for optimal operation Marot et al., (2021); Kelly et al., (2020). Although the ongoing energy transition poses major challenges to grid operators, it concerns other stakeholders as well. Markets must adapt to the decentralized structure while new market players emerge, such as Electric Vehicle (EV) charging operators. Further, with the ongoing digitization and build-up of communication systems, the classical power system transforms into a cyber-physical energy system (CPES) Steinbrink et al., (2018). All these new challenges introduce a new layer of complexity into overall power grid operation, which has to be considered for the success of the energy transition.
Traditionally, grid operation has mostly relied on traditional optimization approaches for Optimal Power Flow (OPF) problems. Due to the non-linear and non-convex nature of OPFs, grid operators often fall back on relaxation techniques or heuristic approaches Srivastava et al., (2023). However, relaxation techniques may yield imprecise results, and cannot ensure the solution’s optimality, thus casting doubt on their effectiveness in application. Additionally, classical optimization approaches struggle to scale to real power grids such that exact results can be achieved in a reasonable time, thus relaxation techniques are applied Molzahn and Hiskens, (2019). Noisy or missing measurements cannot reliably be handled by classical approaches Liu et al., (2020). Therefore, practitioners are are exploring deep-learning solutions Marot et al., (2020); Marot et al., 2022a for problems involving power flows such as grid operation. Such solutions appear to be a promising alternative to classical approaches, addressing the challenges of time criticality, scalability and reliability of results.
Deep Reinforcement Learning (DRL) techniques can identify and utilize under-exploited flexibilities in power grids, often overlooked by traditional methods and human operators Kelly et al., (2020). By learning from interactions with the grid environment, DRL agents can dynamically adjust to changing conditions and unforeseen events, potentially preventing cascading failures and blackouts Marot et al., (2021); Donnot et al., (2017). Furthermore, its capacity to consider long time horizons aligns with the inherently dynamic nature of power grids Viebahn et al., 2022a . However, the development and training of DRL agents require extensive simulations, as direct interaction with the physical grid is impractical. These simulations often need to abstract from reality and rarely use real data, leading to challenges in transferring the solutions back to real-world application Kaspar et al., (2020). Furthermore, the challenge of large combinatorial action spaces in power grids hinders the direct application of DRL Dulac-Arnold et al., (2021), highlighting the need for handcrafted action spaces and other reduction techniques EI Innovation Lab, (2020); Lehna et al., (2022, 2024). Despite these challenges, the potential for RL to transform power grid management is significant, particularly as power systems aim at meeting decarbonization goals and accommodating more renewable energy sources Prostejovsky et al., (2019); Marot et al., 2022c . Rather than replacing human operators, the goal is to provide RL-driven action recommendations for themViebahn et al., 2022a ; Marot et al., 2022b . While proof-of-concept results are impressive, research on DRL for power grid control is still in its early stages and a significant gap remains before deployment can be achieved.
The complex interconnectedness within a power grid can naturally be modelled as a graph, in which nodes and edges correspond to elements and connections in the grids Viebahn et al., 2022a . Since GNNs are specifically designed for such graph-structured data, they are highly suitable for modelling interdependencies in power systemsLiao et al., (2021). Furthermore, they excel at capturing spatial relationships and enable an effective feature extraction from the grid. While feed-forward neural networks struggle to produce accurate results when the grid’s topology, and thus the input dimensionality, changes, GNNs are more robust to modifications of the graph structure. This is an advantage, as several grid actions such as bus-bar splitting can transform a single node in a graph into two distinct nodes (or alternatively, combine two nodes into one). This is uncommon in other types of networks Donon et al., (2020). The distinctive configuration of substations requires tailored methods to account for specific network constraints and leverage network flexibilities. Similarly, efficiently managing long-range dependencies such as non-local effects and the rapid propagation speed of electricity pose unique challenges for power systems. In other words, the effectiveness of GNNs across various topologies and different power grids still requires extensive research Ringsquandl et al., 2021a . Furthermore, the interpolation capabilities of GNNs are instrumental in reconstructing missing information and smoothing out noisy measurements Kuppannagari et al., (2021). This addresses practical issues in power systems where sensors may experience connectivity problems, leading to incomplete or unreliable data.
The combination of GNNs and RL represents a synergy that harnesses the strengths of both paradigms. GNNs provide a powerful tool for feature extraction and representation in graph-structured data, enhancing the RL agent’s understanding of the complex relationships within power grids. As pointed out by Munikoti et al., (2023), the performance of RL agents strongly depends on the state encoding, and GNNs are much better encoders for graph-structured environments. Incorporating them into RL has the potential to enhance the agent’s ability to learn from the graph structure. This has the potential for more informed decision-making, better adaptability to changing network conditions, and improved generalization across diverse scenarios and topologies.
This survey explores recent trends in leveraging GRL for power grids, specifically focusing on transmission and distribution networks as well as other power grid applications such as the energy market, communication networks within power systems, and EV charging management. We focus on DRL models that utilize GNNs to capture the graph-structured state space of power grids while leveraging DRL for sequential decision making.
1.1 Adressed Use Cases
1.1.1 Transmission Grid Control
To ensure safe and reliable transmission, human experts traditionally manage and monitor the power grid in control rooms manually. However, the integration of stochastic and dynamic elements, such as renewable energy sources increases the complexity of power grid operation, starting to surpass human expertise Marot et al., (2021). Therefore, automation and data-driven optimization become necessary. The challenge lies in automatically controlling large-scale power grids, which requires intricate yet dependable decision-making in a large decision space.
Many approaches for transmission grid management focus on controlling electricity generation, or loads, e.g., re-dispatch Kamel et al., (2020); Bai et al., (2023); Fuxjäger et al., (2023) or load shedding Larik et al., (2019). An effective alternative solution can be achieved by managing the power grid through topology control, such as substation switching. This grid reconfiguration allows for efficient rerouting of power flow while using existing infrastructure. This is less costly than generator re-dispatch and minimizes disruptions to the existing system. Despite preliminary studies on grid topology control in literature, the large action spaces, and non-linear nature of the problem hinder the practical deployment of existing methods in real-world scenarios.
Researchers have explored multiple approaches to optimize power flow on the popular Grid2op Donnot, (2020) environments of the Learning to Run a Power Network (L2RPN) challenges Marot et al., (2020). Such a simulation and testbed is crucial for developing methods to tackle congestion and enhance grid reliability in networks at different scales. Therefore, research on RL for topology control so far is dominated by Grid2op, highlighting its pivotal role in advancing the field.
1.1.2 Voltage Control for Distribution Grids
The distribution side of power systems, characterized by the need for decentralized decision-making, presents its own challenges. The integration of distributed generation and ramping up the connection of new loads, such as EVs and heat pumps, poses challenges in voltage control for distribution grids Srivastava et al., (2023). Additionally, more sensors also come with an increased amount of available data, which can improve the training of machine learning models such as GRL. In contrast to heuristics, such approaches can use available data from multi-model sources without major effort. Furthermore, the exploration of more comprehensive optimization problems, which use various voltage control options within distribution grids, such as the integration of novel assets or leveraging the functionalities of distributed energy management systems, are promising in the advancement of distribution grid technology. In this context, RL is particularly suitable as it can deal with multiple objectives at the same time.
1.1.3 Other Use cases
GRL shows potential for related applications such as new energy markets, communication networks for power operation, and cross-sectoral applications like EV charging scheduling with power grid constraints. In energy markets, traditional centralized methods struggle with privacy and scalability, making distributed methods like multi-agent RL necessary. GNNs have the potential to enhance these methods by improving modeling accuracy and scalability. In addition to the physical transmission of power, modern energy systems rely on efficient information transmission for monitoring and control. This requires optimized communication networks to route packets effectively without losing critical data. GRL shows promise to optimize these networks, supporting physical grid operations. Finally, GRL is also promising in addressing challenges from the rapid growth of electromobility by determining charging station recommendations and navigation to address power grid boundaries and constraints, thus aiding the integration of electromobility into the broader energy landscape.
1.2 Contribution and Structure
The main contributions are as follows:
  • We are the first to comprehensively analyze existing GRL approaches for different power grid use cases, categorizing them based on the specific scenarios they address. Our analysis particularly emphasizes applications in distribution and transmission grids. Within the distribution grid, we differentiate between regular voltage control and emergency situations, while for the transmission grid, we concentrate on topology control and the relevant frameworks.
  • We provide an overview of the applied GRL techniques, including states, actions, rewards and analyze the proposed GNN architecture in detail. We identify commonalities and differences between the analyzed methods and point out the most common approaches.
  • We outline the particular advantages of GNNs for RL in power systems, and identify gaps and open problems in existing approaches.
  • We investigate limitations and open challenges of the proposed approach and point out aspects that are crucial for the application of GRL in real-world scenarios.
The papers we analyse in this survey are all published after 2020 up until May 2024, as GRL is a relatively new field. Our selection includes the papers that explicitly address power grids using the combination of GNNs and RL. In particular, we focus on grid operation, which is typically approached for distribution and transmission grids, but we also consider use cases in energy markets, EV applications, and power communication networks if they consider the underlying power grid in their approach.
This review is structured as follows: In Ch. 2, we outline related reviews that deal with methodologies or use cases close to ours. Second, we present the basics of transmission and distribution grids, as well as GNNs and RL in Ch. 3, including the most important concepts used in the analysed approaches. Then, in the main part of this review, we discuss the presented methods in detail and group them by the use case addressed. This part starts with approaches for common problems in transmission grids (Ch. 4) and continues with applications that are typically relevant in distribution grids (Ch. 5). Then, in Ch. 6, we highlight a few relevant papers that cater to different related use cases, such as EV charging. Finally, we give an overall conclusion and outlook.
2 Related Work
In the field of RL as well as GNNs, several more general surveys have been published highlighting different aspects of the respective method or focusing on specific application areas. Broader GNNs surveys that cover the most common GNN architectures are for example presented by Zhou et al., (2020), Thomas et al., (2022) or Wu et al., (2020). Skarding et al., (2021) focus on GNNs for dynamic graphs that evolve over time, while Wu et al., 2022b is an application specific survey that examines GNN approaches for recommender systems. Similarly, in the field of RL, there are several general reviews such as Arulkumaran et al., (2017); Garcıa and Fernández, (2015); Zhu et al., (2023).
Regarding DRL in the context of energy systems, Zhang et al Zhang et al., (2019) provide an overview of existing approaches. They analyze different DRL algorithms that address energy-related problems such as energy management, demand response, electricity market or operational control. There are also more specialized works, such as the review by Vázquez-Canteli et al. Vázquez-Canteli and Nagy, (2019) that examines approaches for RL approaches that address demand response applications in smart grids. However, these works do not include graph reinforcement learning, since it is not yet as well established as RL combined with traditional neural networks and the number of publications has only increased in recent years.
Surveys or literature reviews of approaches combing these two methods are rather rare also in other application areas. Munikoti et al. Munikoti et al., (2023) present a general survey on GRL identifying 80 relevant papers on different use cases. They divide the approaches into two categories: DRL enhancing GNNs and GNNs enhancing DRL. The former deals with DRL applied to architecture search, improving the explainability of GNNs or generating adversarial attacks on GNNs. The latter one focuses on DRL approaches that use GNNs within a DRL setup, which is also the focus of this review. The application areas they investigate are combinatorial optimization, transportation, control of system such wireless networks or material flow.
Fathinezhad et al., (2023) present another survey on GRL that focuses on the basics of GNNs and RL and their relationship. It points out the concepts of how graphs and RL are combined and does not investigate any specific application areas. Furthermore, the authors particularly focus on multi-agent RL, where GNNs are used to enable communication and information exchange between agents. We, in contrast, focus more on the use of GNNs as feature extractors for data that has an inherent graph structure. Fathinezhad et al., (2023) mention only one energy-related application (Pei et al., (2023)), but do not analyze it in detail, since application-specific approaches are not the focus of their survey.
The survey presented by Nie et al., (2023), also examines GRL from a more methodological point of view. They analyze in detail, how RL can be used to improve GNNs and to solve graph problems in general. The survey covers several application fields which are transportation network optimization, E-commerce recommendation systems, structure prediction and generation in medical research, and modelling and controlling of the COVID19-virus.
The survey presented by Mazyavkina et al., (2021) explores GNNs as representation learners for graph structures problem in the RL setup, however the authors concentrate exclusively on on combinatorial optimization problems.
In Mendonca et al., (2019) give a detailed overview of how different graph algorithms can be used to enhance RL by abstracting primitive actions into temporal macro-actions. Similarly, Pateria et al., (2021) analyze hierarchical RL which often involves graph-based approaches, e.g. to discover subtasks, but both works do not consider GNNs in that context.
In all of the above mentioned related works, the application to the energy system is not reviewed. A review focusing on GNNs and their application in power systems is presented by Liao et al., (2021). This review, published in 2022, analyses different application-specific approaches to fault scenarios, time series prediction for renewable energy, data generation and power flow calculation with respect to the GNN architecture used. Liao et al. point out that the reviewed GNN methods typically show better performance than traditional neural networks in the experiments conducted. However, there are still several gaps and open questions that require further investigation, especially due to the novelty of the field. Liao et al. do not consider approaches using RL.
Regarding the different use cases in power grids, there are related works that specialise in specific problems such as voltage control, e.g. Srivastava et al., (2023) or Murray et al., (2021). However, the works considered are mainly traditional optimisation methods and heuristics, since such approaches are still the most dominant in practice. To the best of our knowledge, a detailed overview on the application of GRL in the context of energy systems do not yet exist. Therefore, we aim to fill this gap with this work.
3 Fundamentals: Power Grids, Graph Neural Networks and Reinforcement Learning
3.1 Power Grids
Power grids are essential to modern society and a crucial part of today’s infrastructure. In the face of the energy transition, power system engineers encounter challenges across all aspects of the grid. Its general purpose is to transport electricity from generation units to consumers, which typically are not located in the same area. Traditionally, generation was centralized, e.g., at fossil-fueled or nuclear power plants, resulting in unidirectional power flow from generation to consumption. With the worldwide rise of renewables, generation units are spread across the power system resulting in a decentralized structure and bi-directional power flow.
With the electrification of the traffic and heating sectors, the consumption side of power grids is also undergoing a major shift. Not only does the overall consumption increase with this sector coupling, but also new simultaneity of consumption have to be considered by power system engineers, e.g., simultaneous heating in winter or swing season resulting in a simultaneous heat pump operation over a grid area.
Power grids are typically divided into two levels: transmission and distribution. These levels differ in voltage levels, purpose, and the resulting characteristics, and are split by substations (see 1). We elaborate on both power grid levels in the following subsections.
Figure 1: Visualization of the power grid structure with transmission and distribution level
3.1.1 Transmission grids
Transmission grids and their assets operate at high voltages ranging from 110 kV up to 765 kV. The exact voltage level is country-specific, therefore a clear distinction of voltage levels cannot be made. For example, in Germany the 110 kV network is considered as the high-voltage distribution grid. However, the purpose of transmission grids is across all countries the same. It aims to transport large quantities of electrical power over a large distances. The distances vary from a few hundred kilometers (e.g., connecting off-shore wind in Northern Germany with the West-German industry) to a few thousands (e.g., China’s cross-country interconnections).
Transmission grid operation aims for at minimum N-1 secure operation, i.e., of one asset in the grid fails, the remaining grid is not overloaded. Therefore, transmission grids are typically built in meshed structures with a lot of redundancies installed, e.g, multiple transformers and busbars at substations. The structure of transmission grids results in a highly-complex system where optimization problems have to be solved on a large scale with robust results.
One remedial action to prevent grid congestion, e.g., transformer or line overloading is re-dispatch which refers to the change of generator injections. As generation and consumption in power system have to be balanced at all times, the change of one generator set point has to result in a set point change of another generator. This can result in a step down of renewable generators and a step up of fossil-fueled generators, which is an undesirable action, both in terms of costs and CO2-neutrality. Therefore, other means of flexibility are researched such as topology control of the transmission system. Controlling the switching state in substations, the topology of the grid can be modified and help in decreasing or even completely avoiding the need for re-dispatch. Considering the topology into optimization is a challenge itself, as it results in a Mixed-Integer Non-Linear Problem. Here, AI solutions such as RL can help and research on these challenges are on-going Marot et al., (2020).
3.1.2 Distribution grids
Distribution grids are operated on lower voltage levels, typically lower than 110 kV (e.g., 12.47 kV and 20 kV) until voltage levels of e.g., 120 V and 230 V. The purpose of distribution grids is to deliver the transmitted electrical power from the transmission grids to its individual customers, such as industrial consumers at medium-voltage level and directly to households.
The geographical area covered by distribution grids is smaller than for transmission grids, e.g., one low-voltage grid serving a village with electricity. As the impact of asset failures is rather low in comparison to transmission grids, distribution grids are often not designed considering the N-1 security criterion. Therefore, radial or open-loop structures are common in distribution grid operation.
This results in a higher volatility to voltage fluctuations due to changing generation or consumption over time. With up-ramping installations of photovoltaic generation, this voltage fluctuations can cause voltages to exceed the limits set by the respective DSO. Traditionally, regulated distribution transformers, shunt capacitors, and voltage regulators are used for voltage control Srivastava et al., (2023). With ongoing digitization efforts, the possibilities for voltage control are increasing. Inverter-based technologies, such as smart PV inverters Howlader et al., (2020), vehicle-to-grid Gonzalez Venegas et al., (2021), and stationary batteries Stecca et al., (2020) are flexibilities that become necessary to be tapped by grid operators to maintain the system in a secure state. Due to the high number of distribution grids in the power system, optimized control strategies need scalable and robust solutions, where RL-based solutions can help.
3.1.3 Grid models
In research there are multiple instances of synthetic grid models. Their purpose is to have a common benchmark for researchers across power system engineering to test new concepts and to be able to compare results with each other. There are multiple examples in literature for benchmark grid models Babaeinejadsarookolaee et al., (2021); Meinecke et al., 2020a ; Meinecke et al., 2020b ; Strunz et al., (2014). In the above mentioned use cases of transmission and distribution system control, the IEEE test cases are the most commonly used grid benchmarks, available in multiple power system calculation tools like MATPOWER Zimmerman et al., (2011), pandapower Thurner et al., (2018), and DIgSILENT PowerFactory DIgSILENT GmbH, (2024).
However, these benchmark cases are simplified cases of power systems which do not fully capture the complexity and challenges of power system operation. These grid models are mostly simple bus branch models with loads and generators connected to the buses. More sophisticated assets like, inverter-based generator controls, transformer tap changers, shunt elements, etc., are not common for such grid models but are common practice in real-life power system operation. Therefore, methods benchmarked on such synthetic grid models have to be treated cautiously as the application in practice requires further considerations.
3.2 Graph Neural Networks
Figure 2: Left: Visualization of the general message passing scheme in GNNs - The target node (orange) receive messages $m_{u}i$ from its neighbours and aggregates them. The messages can be constructed from information of both, the target and neighbouring node depending on the message passing scheme. Right: Illustration of a GNN - The graph is input to the GNN layers which compute node embeddings based on the messages from neighbouring nodes. As indicated in orange, this is done for each node in the graph. After all embeddings are computed, an activation function is applied. This is repeated for a given number of layers. In the end, the GNN outputs a graph with new node feature from which prediction can be made.
GNNs are designed to extract information from graph-structured data by applying multiple layers of graph convolutions. They can be interpreted as a generalisation of Convolutional Neural Networkss to non-euclidean structured data. The general idea is to combine information from local regions of the inputs in a learnable way and to grow these local regions from layer to layer. In this way, CNNs learn increasingly abstract features from the input data from layer to layer. CNNs perform extremely well on grid-structured data, such as images. Here, the neighbourhood of a pixel consists of the neighbouring pixels, forming a regular grid. However, many real-world phenomena involve relationships or complex dependencies that cannot be represented as regular grid structures without losing information. For example, in a regular grid every node has the same number of neighbours, but in power grids not every components is connected to the same number of power lines.
Graphs consist of an unordered set of nodes and edges, where the edges define the neighbourhood of a node. Graphs can therefore be used to model complex relationships, such as neighbourhoods of arbitrary size or multiple types of edges or nodes. They can have attributes that describe properties of nodes and edges, such as node features or the strength of an edge. This makes graphs a perfect model for many real-world applications, such as power grids. In addition, feed-forward neural networks working on Euclidean data typically treat nodes as independent samples. This means, they neglect the relationships between nodes, or stack them unsystematically into a vector. GNNs, on the contrary, make use of the information about node connectivity. They can solve all common learning tasks, i.e. classification, regression, and clustering, for entire graphs, and at node- or edge level.
3.2.1 Message Passing
GNNs update the embeddings of the graph nodes by repeatedly aggregating information of their neighbourhoods in a learnable way. This general scheme is called message passing as each node updates its embedding based on the messages it receives from its neighbours. The representation $\bm{h}_{u}$ of a target node $u$ generated by a general message passing layer is computed as:
$\bm{h}_{u}=\sigma\left(\bm{x}_{u},\bigoplus\limits_{v\in\mathcal{N}(u)}\psi(\bm{x}_{u},\bm{x}_{v})\right),$ (1)
where $\psi$ corresponds to a message function equipped with learnable weights that compute a message between node $u$ and its neighbour $v$. $\bm{x}_{u}$ and $\bm{x}_{v}$ are the respective node embeddings. $\mathcal{N}(u)$ is the neighbourhood of node $u$, $\bigoplus$ refers to an aggregation function that defines how messages are passed Bronstein et al., (2021) and $\sigma$ is an activation function. There exist various different implementations of message passing layers, the scheme shown above, being the most general one Bronstein et al., (2021).
Spatial Graph Convolution
A simple form of message passing is the spatial graph convolution. Here, messages are the features of the neighbouring nodes transformed using a learnable weight matrix. The aggregation corresponds to the summation operation. Morris et al., (2019) for example propose such an intuitive formulation:
$\bm{h}_{u}^{\prime}=\sigma(\bm{W}_{u}\bm{h}_{u}+\sum_{v\in\mathcal{N}(u)}\bm{W}_{v}\bm{h}_{v})$ (2)
with $\bm{W}_{u}$ and $\bm{W}_{v}$ being learnable weights matrices, also referred to as filters. Typically, the weights are shared across all nodes and neighbours. This follows the concept of parameter sharing in CNNs.
Graph Attention Network (GAT)
A method commonly used to improve the performance of a GNN model is to equip the graph convolution with an attention mechanism. Such attention layers are a special case of general message passing Bronstein et al., (2021). Hereby, attention coefficients are learned for each connected pair of nodes. They are computed from the features of the neighbouring nodes and the target node and determine the influence a neighbour has on the target node. Veličković et al., (2018) give a definition of such an attention convolution that would extend Eq. 2 to:
$\bm{h}_{u}^{\prime}=\sigma(\bm{W}_{u}\bm{h}_{u}+\sum_{v\in\mathcal{N}(u)}\alpha_{u,v}\bm{W}_{v}\bm{h}_{v})$ (3)
where $\alpha_{u,v}$ refers to the attention coefficient for node $v$ in the neighbourhood aggregation of node $u$ which indicates the importance of node $u$ to node $v$. It is computed as:
$\bm{\alpha}_{u,v}:=\text{softmax}_{v}\left(\sigma(\bm{a}^{T}[\bm{W}_{a}\bm{h}_{u}||\bm{W}_{a}\bm{h}_{v}]\right)$ (4)
with $\bm{a}$ and $\bm{W}_{a}$ being weight vector and matrix respectively. $\sigma$ again refers to an activation function, Veličković et al., (2018) for example use ReLu. The ||||| | corresponds to the concatenation operation, so the transformed features of both nodes are concatenated before the attention mechanism $\bm{a}$ is applied to produce the attention coefficients. The coefficients are normalized using softmax to make them comparable across the neighbours of the target node.
Spectral Graph Convolution
[…]
Here, $\bm{U}$ is the matrix of eigenvectors of the normalized graph Laplacian $\bm{L}=\bm{I}-\bm{D}^{-\frac{1}{2}}\bm{A}\bm{D}^{-\frac{1}{2}}$ and is determined by the eigendecomposition $\bm{L}=\bm{U}\bm{\Lambda}\bm{U}^{T}$. Convolution in the spectral domain then is simply an element-wise multiplication, after which, the convolved signal is transformed back into the graph domain:
[…]
Where $\bm{U}^{T}\bm{g}$ is the filter in the spectral domain. Since the normalized graph Laplacian $\bm{L}$ is composed of the degree matrix $\bm{D}$ and adjacency matrix $\bm{A}$ ($\bm{I}$ being the unit matrix), intuitively its eigenvectors and eigenvalues indicate the major directions of information diffusion through the graph.
The spatial and spectral formulations of GNNs are equivalent, but due to the high computational cost of spectral GNNs resulting from the eigendecomposition at each step, spatial GNNs are more commonly used in many practical applications. In physical systems such as power grids, however, they appear relatively popular.
GNN Training
Since GNNs are differentiable functions, they can be trained just like ordinary NNs, i.e., using gradient descent, backpropagation, batches or mini-batches. Commonly used loss functions include the negative log-likelihood of softmax functions for node or graph level classification. For link prediction, pairwise node embedding losses such as cross entropy or Bayesian personalised ranking loss are common.
Challenges
Due to their specific functionality, GNNs suffer from oversmoothing, oversquashing and scalability problems. Oversmoothing describes the problem that with an increasing number of layers, the features in each node become increasingly similar. This problem can for instance be addressed by regularisation or normalisation. Oversquashing refers to the distortion of information from distant nodes and is difficult to handle Giovanni et al., (2024). Regarding scalability, a GNN can either train on the entire graph and thus keep the full graph in memory. This may not be feasible for very large graphs. Or it can process mini-batches in the form of subgraphs, which leads to exponentially growing computational complexity with respect to the number of GNN layers Ding et al., (2022). While many ways to tackle this problem have been proposed, most of them sampling-based Wu et al., 2022a , scalability still remains a challenge.
3.2.2 Other Architectures used in the analyzed Approaches
GraphSage
The architecture proposed by Hamilton et al., 2017a is a special case of spatial graph convolution based on sampling. Instead of aggregating the entire neighbourhood of a node in each layer, a fixed number of neighbours of the target node are randomly sampled. The neighbours are then aggregated using a permutation invariant function such as mean, max or sum. They propose a loss function that allows unsupervised training of GraphSage. It consists of two terms, one enforcing that nodes that are close in the input graph have similar embeddings and the other pushing apart the embeddings of two nodes that are far apart in the graph.
Graph Capsule Networks
The idea of Graph Capsule Networks proposed by Verma and Zhang, (2018) is to capture more informative local and global features. This is done by constructing a capsule vector that contains enough discriminative features to allow proper reconstruction. These vectors are constructed using a capsule function that maps the input features of neighbouring nodes to higher order statistics. In a simple form, this could mean that the resulting vector contains the neighbourhood’s mean, the standard deviation and statistics of higher orders, depending on the given dimensionality of the capsule vector. The second key component of graph capsule networks is the aggregation function which is based on the covariance of the graph as it provides information such as norms or angles between node features.
3.3 Reinforcement Learning
RL is a key branch of machine learning thatfocuses on training agents to make sequential decisions in dynamic environments to maximize cumulative rewards. It primarily uses the framework of Markov Decision Processs to model decision-making problems. An MDP is defined by the tuple $M=(S,A,R,T,\gamma,H)$, capturing essential elements of the agent’s interaction with its environment. Figure 3 shows the schematic interaction between agent and environment in the RL framework.
Figure 3: The agent-environment interaction is a cyclical process where the agent selects actions based on the current state, leading to state transitions and rewards, guided by a policy $\pi$, hence generating a sequence of states, actions, and rewards.
In this formalism, $S$ is the state space of all possible states, and $A$ is the action space of feasible actions. The reward function $R$ maps states and actions to real-valued rewards, providing immediate feedback. The transition function $T$ describes state transitions in response to actions. The discount factor $\gamma$ balances the importance of future rewards against immediate ones. Finally, the horizon $H$ defines the length of an episode, consisting of a sequence of states, actions, and rewards. In contrast, partially observed MDPs incorporates situations where the agent has incomplete knowledge about the current state. They provide a more realistic framework for many real-world problems where the agent must make decisions based on partial and uncertain information about the environment.
The primary objective of an agent is to learn a policy function $\pi(s)$ that prescribes actions to states, aiming to maximize the expected cumulative discounted sum of rewards over the time horizon $H$ that defines the length of the episode. Here, $\pi^{*}$ denotes the optimal policy.
$\pi^{*}=\arg\max_{\pi}\mathop{\mathbb{E}}\left[\sum_{t=0}^{H}\gamma^{t}r(s_{t},a_{t})\right]$ (7)
RL algorithms are commonly classified into two main types: model-free and model-based methods. Model-free methods operate without requiring knowledge of the environment’s transition functions; instead, they utilize the experiences gathered by the agent. These methods can be subdivided into two primary categories: policy-based and value-based methods, depending on their approach to solving an MDP. In contrast, model-based approaches focus on scenarios where the transition function is either known or can be learned. Examples of model-based methods include Monte Carlo Tree Search (MCTS) algorithms like AlphaZero Silver et al., (2016), MuZero Schrittwieser et al., (2020) and EfficientZero Ye et al., (2021). In the following, we introduce two important concepts for model-free RL, namely value-based and policy-based learning, as well as actor-critic approaches. Then, we will briefly present the widely used model-based technique MCTS.
3.3.1 Model-free RL
Value-based learning
Value-based learning estimates the value of state-action pairs to select actions that maximize these values, using the value function $V^{\pi}(s)$ to represent the expected cumulative reward of following policy $\pi$ from state $s$.
$V^{\pi}(s)=\mathop{\mathbb{E}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t})\mid\pi,s_{0}=s\right]$ (8)
Similarly, the action-value function $Q^{\pi}(s,a)$ represents the expected sum of future discounted rewards, beginning from state $s$, executing action $a$, and subsequently adhering to a given policy $\pi.$
$Q^{\pi}(s,a)=\mathop{\mathbb{E}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid\pi,s_{0}=s,a_{0}=a\right]$ (9)
The value function has a key recursive property linking the value of state $s$ to the values of subsequent states $s^{\prime}$, which is fundamental to many value-based RL techniques. This is expressed by the Bellman equation
$Q^{\pi}(s,a)=r(s,a)+\gamma\sum_{s^{\prime}}p(s,a,s^{\prime})\max_{a^{\prime}}Q^{\pi}(s^{\prime},a^{\prime})$ (10)
where $p(s,a,s^{\prime})$ models the state transition dynamics. In value-based approaches, finding the optimal policy involves identifying the optimal value function $Q^{*}(s,a)=\max_{\pi}Q^{\pi}(s,a)$. Explicit solutions to the Bellman equation are possible only when the dynamics function is known Sutton et al., (1999), so approximation methods are typically used. In the following, we present two such approximation methods.
  • Q-Learning is an off-policy learning method Watkins and Dayan, (1992) that aims to derive an optimal policy by directly updating values in a Q-table, a lookup table where each entry $Q(s,a)$ estimates the expected cumulative reward for taking action $a$ in state $s$. Q-Learning approximates the optimal action value function $Q^{*}$ through the following iterative updates.
    $Q(s,a)\leftarrow Q(s,a)+\alpha\left[r(s,a)+\gamma\max_{a}Q(s^{\prime},a)-Q(s,a)\right]$ (11)
    Here, the agent explores the environment with a behavior policy, updating the Q-table based on observed rewards and state transitions.
  • Deep Q-Networks use neural networks to approximate action-value functions in high-dimensional input spaces, minimizing the mean-squared error between current and target Q-values Mnih et al., (2015). It introduces techniques like replay buffers and target networks to enhance training stability. Variants such as Double Deep Q-Network (DDQN) Hasselt et al., (2016), dueling DQN Wang et al., (2016), and Rainbow Hessel et al., (2017) further improve performance by addressing overestimation and efficiency issues.
Policy-based learning
Policy-based learning directly estimates policies without intermediate value functions. It optimizes parameterized policies $\pi_{\theta}(a\mid s,\theta)$ that specifies the probability of action $a$ given state $s$ and parameters $\theta$ in order to maximize expected cumulative rewards. Unlike value-based approaches, policy-based methods update parameters using gradient-based optimization and are suitable in managing continuous action spaces and stochastic policies Sutton and Barto, (2018).
The objective function $J(\theta)$ aims to maximize the true value function $v_{\pi_{\theta}}(s_{0})$ from the initial state $s_{0}$. According to the policy gradient theorem Sutton et al., (1999), $J(\theta)$ is proportional to the sum of the action-value function multiplied by the gradient of the policy:
$J(\theta)\propto\sum_{s}\mu(s)\sum_{a}q_{\pi}(s,a)\nabla\pi_{\theta}(a\mid s,\theta)$ (12)
Here, $\mu(s)$ is the distribution under $\pi$, $q_{\pi}(s,a)$ is the action-value function, and $\nabla\pi_{\theta}(a\mid s,\theta)$ is the gradient of $\pi$ with respect to $\theta$. This gradient is shown below and can be computed as the product of the policy probability and the gradient of the logarithm of the policy.
$\nabla_{\theta}\pi_{\theta}(a\mid s,\theta)=\pi(s,a)\nabla_{\theta}\log\pi_{\theta}(s,a)$ (13)
where $\nabla_{\theta}\log\pi_{\theta}(s,a)$ is the scoring function. Thus, the update of the policy parameters proceeds in the direction of the gradient of the objective function: $\Delta\theta=\alpha\nabla_{\theta}J(\theta)$, where $\alpha$ is the learning rate.
Actor-Critic methods
Actor-Critic algorithms combine the strengths of value-based and policy gradient-based learning. The actor learns policies to maximize rewards, while the critic evaluates these policies by estimating the value function. This framework addresses the limitations of both approaches and is fundamental to various RL algorithms, including Asynchronous Advantage Actor-Critic (A3C) Mnih et al., (2016), Deep Deterministic Policy Gradient (DDPG) Lillicrap et al., (2015), Proximal Policy Optimization (PPO) Schulman et al., (2017), and Soft Actor-Critic (SAC) Haarnoja et al., (2018).
For instance, A3C updates policy and value networks asynchronously across multiple threads, using an advantage function to guide actions and entropy regularization to enhance exploration. Similarly, DDPG, tailored for continuous action spaces, simultaneously learns a state-action value function (critic) and a policy (actor), employing experience replay and target networks for stable learning. In addition, PPO uses a clipped surrogate objective for smooth policy updates, balancing exploration and exploitation, which makes it a popular choice in RL research. Furthermore, SAC combines actor-critic methods with entropy regularization, training a policy network and two Q-value networks concurrently to encourage diverse action exploration and reduce overestimation.
While established algorithms such as PPO have been extensively used, newer innovations are often overlooked, particularly in power grid control. For instance, Bigger, Better, Faster (BBF) Schwarzer et al., (2023) is an advanced method that addresses the challenge of scaling networks effectively with limited samples. It employs a 15-layer ResNet architecture with widened layers, a high replay ratio Fedus et al., (2020) with periodic network resets, and adaptive strategies like dynamic update horizon and discount factor schedules. BBF discards NoisyNets Fortunato et al., (2017) in favor of weight decay for regularization, and outperforms other state-of-the-art agents in both computational efficiency and task performance, making significant strides in optimizing DRL for constrained data environments.
3.3.2 Model-based RL
Monte Carlo Tree Search (MCTS)
Both value-based and policy-based approaches in RL operate as model-free methods, meaning they do not utilize the model of the environment to plan ahead by simulating future steps. This is where MCTS Browne et al., (2012) comes into play, it is a heuristic search the combines the accuracy of tree search with the power of random sampling to efficiently explore large state spaces. The algorithm builds a search tree incrementally, where nodes represent states and edges represent actions.
The process begins with selection, where the algorithm chooses the most promising child nodes from the root until it reaches a leaf node. If the leaf node is not terminal, the expansion phase adds one or more child nodes. Next, during the simulation phase, the algorithm runs a simulation from these new nodes to a terminal state, typically using random actions. Finally, in backpropagation, the results of the simulation are used to update the values of the nodes in the path from the leaf to the root, propagating the success or failure of the simulation.
MCTS effectively balances exploring new actions and exploiting known high-reward actions by using the Upper Confidence Bound for Trees (UCT) formula to select nodes. This balance has made MCTS a powerful tool, with notable implementations such as AlphaZero Silver et al., (2016), MuZero Schrittwieser et al., (2020), and EfficientZero Ye et al., (2021) achieving superhuman performance in complex games.
AlphaZero, designed to master games like chess, shogi, and Go, utilizes deep neural networks combined with MCTS and relies on predefined game rules to achieve superhuman performance through self-play and RL Silver et al., (2016). MuZero extends this approach by generalizing to environments where the rules are not known in advance, integrating RL, MCTS, and learned models to predict environmental dynamics Schrittwieser et al., (2020). EfficientZero builds on MuZero and achieves super-human performance on the 100k Atari benchmark, surpassing previous state-of-the-art results significantly Ye et al., (2021). It introduces innovations like self-supervised consistency losses for accurate next-state predictions and end-to-end value prefix prediction to handle state aliasing issues. These enhancements improve exploration and action search capabilities, making EfficientZero highly effective in data-limited scenarios.
4 Graph Reinforcement Learning for Transmission Networks
Management and control are crucial issues in electricity transmission, requiring grid operators to meet all demand at all times. This task is becoming increasingly difficult due to rising demand and the variability from more renewable energy sources. Consequently, grid operation is shifting to adapt to generated power rather than predicted demand Viebahn et al., 2022b . Previously, grids were load-driven, adjusting generation based on estimated demand, but now traditional control methods are less effective as optimization problems grow more complex.
Traditional optimization methods face scalability issues with non-linear power systems, prompting a move towards more adaptive strategies. DRL is emerging as a possible solution for managing uncertainties and constraints in transmission grids Marot et al., (2021).
New grid operation approaches can’t be tested in real scenarios, so they are validated through simulations. Grid2Op Donnot, (2020) is the most commonly used environment, simulating grid operations and evaluating DRL agent performance across various scenarios. This environment is used in the Learning to Run a Power Network (L2RPN) Marot et al., (2020) challenges, which encourage the development of RL methods for grid monitoring and control.
Grid2Op Donnot, (2020) provides several actions, such as generator redispatch to balance supply and demand and curtailment of renewables to regulate output. A significant feature of the L2RPN challenges is real-time control of grid topology using bus switches and busbar splits Silver et al., (2017). This can avoid renewable energy curtailment and improve power transmission efficiency. Topological measures offer a low-cost and dynamic response to congestion and overload.
By training in realistic simulation environments, RL agents can enhance grid efficiency and resilience. However, the vast state and action spaces in transmission grids are challenging. Promising approaches to reduce action space involve heuristic methods Lehna et al., (2023); EI Innovation Lab, (2020); Chauhan et al., (2022) and hierarchical strategies that separate high-level topology decisions from specific actions Manczak et al., (2023); van der Sar et al., (2023); Yoon et al., (2021); Liu et al., (2024).
DRL solutions from these challenges have developed effective control strategies leveraging topological flexibility EI Innovation Lab, (2020); Lehna et al., (2023). Few approaches currently use GNNs, which are advantageous for topological actions by directly learning from the grid’s graph structure. Therefore, in this chapter we analyse existing GRL approaches. The success of DRL solutions in these competitions demonstrates their potential in power system operations. However, simulations can’t cover all aspects of real grid operations, and Grid2Op abstracts from reality by using standardized IEEE grids, potentially limiting real-world applicability. This focus may overlook unique issues of actual power grids.
We identified eight GRL approaches that address grid operation in transmission networks. Their primary objective is to maintain network stability under dynamic conditions in real-time. Thanks to a well-organized pre-dispatch schedule and the power system’s robustness, specific actions are typically unnecessary in routine situations. Consequently, all approaches activate agents only in critical states, such as heavily loaded lines. When no lines are heavily loaded, a "do-nothing" action is implemented, and the agent intervenes only when loading surpasses a threshold. This control mechanism prevents overacting and minimizes costs, as actions are associated with costs and should not occur when the grid is stable. This feature is common among most RL approaches EI Innovation Lab, (2020); Lehna et al., (2023) and GRL approaches Xu et al., (2020); Taha et al., (2022); Yoon et al., (2021); Sar et al., (2023). For example, Sar et al., (2023) employs a hierarchical RL strategy where the highest level agent is a rule-based mechanism triggered by critical line loading.
Table 1: Overview of GRL approaches proposed for transmission grids. We categorize approaches by the used RL algorithm and GNNs, specifying action types and emphasizing topological actions alongside others. Grid sizes indicate the number of buses evaluated. The focus column describes distinctive aspects of each approach.
env Model Control Algorithm Action GNN Grid Size Focus
grid2op Xu et al., (2020) Double-Q-Network Topo GCN IEEE 14 simulation-constrained RL
Taha et al., (2022) MCTS for acion selection Topo GCN with residual connections IEEE 118 GCNs for power flow estimation
Xu et al., 2022a Double-Dueling Q-Networks Topo & redispatch GAT IEEE 118 MCTS for action space reduction, multiple agents with different sub action spaces
Sar et al., (2023) Multi-Agent ISACD, DSACD & PPO Topo GCN IEEE 5 Hierarchical RL
Yoon et al., (2021) SMAAC Topo GNN Transformer IEEE 5, 14, 118 Afterstate representation, goal topology actions
Zhao et al., (2022) PPO redispatch, curtailment, battery storage GraphSAGE IEEE 118 Representation of Power Grids, simulation of bus additions
Qiu et al., (2022) SMAAC Topo GAT IEEE 5,14, 118 (subset) Afterstate representation, attention mechanism
other Wu et al., 2023b Primal-Dual Constrained TD3 Active and reactive power control, Battery operation Cplx-STGCN IEEE 14, 30 Feasible control for SDOPF optimization
4.1 RL Framework
Reward
When an action other than "do-nothing" is required, the agent is activated. The goal of the approaches analyzed here is to manage line flows, mitigate congestion, and minimize costs. The reward for GRL agents is based on these aspects. For example, Zhao et al., (2022) and Wu et al., 2023b combine operational costs like power loss and generator dispatch with penalties for constraint violations, such as voltage limits, battery state-of-charge (SoC) constraints, or line flow limits. Xu et al., (2020) use a reward considering line overflow proximity, a gameplay reward reflecting the successful completion of a scenario, and a reward for operational loss. Yoon et al., (2021), Qiu et al., (2022), and Sar et al., (2023) focus on grid efficiency, using the ratio of total served load to total production as a reward. This aims to reduce energy loss from congestion. In contrast, Taha et al., (2022) penalize line overloads with a binary reward counting time steps without overload. Similarly, Xu et al., 2022a aim to improve grid flexibility by penalizing heavily-loaded lines with appropriate weights, targeting potential transmission bottlenecks.
Actions
In critical states, agents can choose to initiate redispatch or alter the grid’s topology by changing bus configurations or line connectivity. Often, lines are reconnected after being disconnected. Since bus-bar splitting by grid operators is cheaper than redispatch actions, it is preferred when possible. However, the action space for bus configurations is very large. These approaches are typically trained on the grid2op environment, which features double-busbar systems. This allows each of the $n$ buses to be set to one of two busbars, resulting in an exponential number of configurations per substation. Simulating each configuration is impractical, especially for larger grids. Even for the small IEEE 14-bus grid, 178 valid actions are available. Approaches like Xu et al., (2020) and Qiu et al., (2022) consider only do-nothing and topological actions. Grid2op restricts bus-bar splitting to one substation per time step, in line with L2RPN challenges and real grid operations. The specific actions considered by each analyzed approach are detailed in Tab 1.
States
Most approaches use the information available in the grid2op environment, though not all use every available feature. For example, Xu et al., (2020) include grid topology, connected elements, and features of generators, loads, and power lines. This data includes the bus-bar index, active and reactive production and consumption, voltage set points and magnitudes, and the active and reactive flow for each power line. Additionally, they use the ratio ($\rho$) between current flow and thermal limits. While state information is generally consistent across GRL approaches, grid size varies by study, affecting the state size. Only Wu et al., 2023b , operating in a different environment, use just voltages in their state representation. Typically, states are modeled as graphs and embedded using GNN approaches, which are described in subsection 4.3 along with the corresponding GNNs.
4.2 Overall Approach and RL Algorithms
While the states, actions, and rewards of the analyzed GRL frameworks are similar, they use different RL algorithms and focus on various problem aspects. Xu et al., (2020), using Double-Q-learning, note that conventional RL agents often violate constraints during exploration. To mitigate this, they employ a soft constraint, replacing actions that violate basic constraints with "do-nothing". During exploitation, they employ an adaptive strategy to select the top $N$ actions with the highest Q-values, checked by the Grid2op simulation to avoid catastrophic results. This strategy improves action selection during early training and saves time when the agent is well-trained.
Instead of Grid2op simulations, Taha et al., (2022) use a GNN to predict the lines loading $\rho$ for different topologies. They estimate state evolution under various actions and select the trajectory maximizing cumulative rewards using MCTS. Starting with the root as the initial state and edges representing actions leading to subsequent nodes, a tree is built iteratively, keeping only nodes with low loads. The above-mentioned GNN predicts the line loading under the corresponding action. After simulating steps with a do-nothing agent, they select the optimal actions by maximizing node values and potential future actions.
Another MCTS-based approach is proposed by Xu et al., 2022a . Their tree structure resembles that of Taha et al., (2022), with leaf nodes representing subsequent overload states from feasible actions. Like Taha et al., (2022), they determine the best action based on the path yielding the highest value. They employ double-dueling Q-networks, training multiple parallel sub-agents to select actions from distinct sub-action spaces derived from MCTS. They use a long-short term strategy to balance immediate and long-term benefits and manage sub-agents effectively. Each agent simulates $n$ actions at overload states, selecting the best action through efficient comparison. They also introduce a constraint to limit topological actions and keep redispatch within physically feasible limits.
In contrast to other approaches, Yoon et al., (2021) focus less on extensive simulation and instead develop an afterstate representation that captures the deterministic grid topology after applying a topological action. This approach directs their RL algorithm towards understanding the stochastic dynamics following each action. They utilize a hierarchical policy with a high-level policy generating desired topologies and a low-level policy executing topology changes. This strategy avoids the need to learn individual actions by focusing on appropriate topologies for critical situations. They explore rule-based approaches, such as CAPA, to efficiently sequence actions, prioritizing substations with high-capacity usage and ensuring timely responses. Their actor-critic algorithm uses the afterstate representation and goal topology predictions to enhance exploration and improve value function determination. Similarly, Qiu et al., (2022) adopt an approach resembling Yoon et al., (2021), employing a different attention mechanism in their Graph Convolutional Network (GCN).
In a novel approach, Sar et al., (2023) employ a 3-level agent system to manage complex decision-making in transmission grids. The highest-level agent assesses the need for action at each time step, often opting for inaction in safe conditions. It uses a rule-based method to identify critically loaded lines. In unsafe scenarios, this agent activates the mid-level agent, which prioritizes substations needing action based on CAPA Yoon et al., (2021), focusing on substations with high line loads. At the lowest level, substation-specific agents manage individual substations by selecting bus assignments within a predefined action space, excluding symmetric actions. This approach stands out for its use of multiple substation-specific agents and the comparison of parameter-shared critics versus independent agents in PPO and SAC.
Unlike other approaches focusing on topological changes, Zhao et al., (2022) tackle the variability in power system topologies caused by factors like extreme weather events and maintenance plans. They emphasize stability and resilience through re-dispatching, curtailment, and battery storage actions rather than altering the grid’s structure. Similar to EI Innovation Lab, (2020), they employ imitation learning to pre-train a PPO RL agent, using a GNN based on GraphSAGE Hamilton et al., 2017b to handle dynamic power network topologies. This approach demonstrates adaptability to topology changes due to emergencies, maintenance, and new bus constructions.
The approach by Wu et al., 2023b tackles a broad problem involving stochastic dynamic OPF with Renewable Energy Sources (RES) and decentralized energy systems, which present significant computational challenges. Their method employs an actor-critic framework where the actor, updated based on the critic’s assessment, anticipates future actions to ensure viability of current decisions. Separate neural networks predict voltages, and critic networks are refined using temporal difference learning. The framework integrates problem constraints using Lagrangian multipliers, leveraging the duality principle to optimize both primal and dual variables through gradient-based primal dual updates. This concept is also applied in similar constrained RL problems like Yan et al., (2023).
4.3 Graph Embeddings
Looking at the representations of the states in these GRL approaches, the grid state is represented as a graph to encompass its elements. Utilizing GNNs, these methods extract features to learn a latent representation of the grid incorporating both its features and topology. While the overall procedure remains consistent across approaches, variations exist in how the power grid is modeled and in the specific GNN architectures used for feature extraction.
Grid2opDonnot, (2020) provides a graph representation where each load, generator, and endpoint of a transmission line is a node. Nodes are connected based on busbar assignments within substations. Each transmission line is represented by two (connected) nodes to account for potential connections to different busbars. This approach simplifies scenarios where substations are split, ensuring consistent handling across different configurations. Node features vary slightly across approaches, typically including all available variables from grid2op, with missing features padded with zeros as necessary.
Xu et al., (2020) utilize the grid2op graph representation and employ a spectral GNN that uses spectral graph convolution to process information within this grid topology. This spectral GNN effectively extracts features from irregular non-Euclidean data. On the other hand, Taha et al., (2022) employ two adjacency matrices to indicate connections at both ends of each line. Each power line is associated with a feature vector that aggregates injections from nodes connected to the same bus at each end. Their processing GNN predicts system dynamics using a structure comprising 2 fully connected layers and 3 graph convolutional layers with residual connections. Following this, two additional fully connected layers make the final predictions. Notably, Taha et al., (2022) use GNNs to predict relative line loads ($\rho$) for an ad-hoc MCTS method. They incorporate a critical step where the GNN is trained on topologies resembling a reference topology. When the GNN’s generalization error increases, they revert the grid’s topology back to this reference state, which has shown positive effects in maintaining grid stability, as affirmed by Lehna et al., (2023, 2024).
Conversely, approaches like Xu et al., (2020); Yoon et al., (2021) employ GNNs directly for learning policy and value functions in DRL contexts. For instance, Yoon et al., (2021) utilize a GNN that comprises transformer-based attention layers for node embeddings, shared across actor and critic heads. Actor heads output mean and standard deviation parameters for action sampling from a normal distribution. Critics similarly process node embeddings to derive scalar values representing value functions. They focus on controllable substations, typically those with more than two elements, to simplify the goal space dimensionality.
In another study on smaller grids like IEEE 5, Sar et al., (2023) utilize three GNN blocks for shared layers in both single and multi-agent settings. The actor incorporates three GNN blocks, while the critic employs one. Although the exact input graph structure isn’t detailed, Grid2op’s typical graph model is assumed.
Meanwhile, Zhao et al., (2022) address dynamic topologies using GraphSAGE within PPO networks to adapt to evolving grid characteristics. GraphSAGE’s unsupervised learning generates node embeddings independently from PPO losses, facilitating generalization to unseen nodes via aggregator functions. It leverages neighborhood sampling, such as random walks, to generate embeddings in changing topologies with varying connectivities and node counts. Their PPO agent integrates GraphSAGE-trained networks with fully connected layers, with shared components between actor and critic, differing only in output for actions and state values.
On a different note, Xu et al., 2022a integrate GATs including multiple attention heads to process power grid states, enabling better adaptation to unpredictable topological changes and improve generalization across different structures. The GAT layers serve as models for both the actor and critic component of PPO, followed by Dense layers to output discrete actions for each agent.
Wu et al., 2023b utilize a Chebyshev Spatio-Temporal Graph Convolutional Network (STGCN) to model dynamics over graph-structured data, specifically tailored for analyzing power grid topologies. The architecture combines temporal convolutional layers with spatial convolutional layers to capture temporal dependencies and spatial interactions within the grid. Temporal convolutional layers employ 1-D CNNs with specified kernel widths and output channels to capture temporal evolution, while spatial convolutional layers extract spatial features using complex-valued transfer functions. This integrated approach yields multiple output channels at each layer.
4.4 Experiments and Evaluation
Most approaches train and evaluate using grid2op environments, since they offer extensive power grid simulations with real data. The analyzed works consider various power grid sizes: Xu et al., (2020) focuses on an IEEE 14-bus system with 20 transmission lines, 6 generators, and 11 loads, spanning 1004 scenarios over 4 weeks at 5-minute intervals. Taha et al., (2022), Xu et al., 2022a , and Zhao et al., (2022) employ the larger IEEE 118 grid. Taha et al., (2022) uses the "l2rpn_neurips_2020_track2_small" environment with adjustments, while Sar et al., (2023) uses the IEEE 5 grid for a hierarchical multi-agent proof of concept. Yoon et al., (2021) tackles multiple sizes including IEEE 5, IEEE 14, and an augmented version of IEEE 118 from L2RPN WCCI 2020. Wu et al., 2023b applies their method on IEEE 14 and IEEE 30 bus systems with wind power data for power flow control using Battery Energy Storage Systems (BESS).
In terms of evaluation, the presented approaches on grid control are not as comparable as one could hope, considering that most are based on the same framework and utilize similar data.
Xu et al., (2020) evaluate their approach using a simulation-constraint double dueling DQN agent compared to a basic double dueling DQN agent on an IEEE 14-bus system. The simulation-constraint agent outperforms the basic agent, maintaining grid stability for longer durations per episode. They also assess the effectiveness of graph convolution by training agents with and without GNN layers, demonstrating that the GNN-enabled agent achieves superior control capabilities. While the current results show promise, further assessment on larger-scale grids is necessary to validate the approach’s applicability and effectiveness.
In contrast, Taha et al., (2022) trained on the larger IEEE-118 grid. The GNN input features represent power lines, with a reduced injection horizon for speed improvement. To gauge topology generalization, the GNN is tested on topologies differing by actions from the reference, revealing RMSE evolution logarithmically increasing with action distance, stable up to approximately 7 actions. Although comparison with feedforward neural networks for line load prediction is absent, their interpretable MCTS agent reduces failure rates significantly, from 15.1% to 1.5%, demonstrating the effectiveness of combining GNNs for prediction and MCTS for action optimization, setting the stage for model-based RL using GNNs as model estimators.
Yoon et al., (2021) validated their Semi-Markov Afterstate Actor-Critic (SMAAC) approach against various baselines, including DDQN and SAC, which performed poorly on medium and large grids due to inefficiencies in exploring all possible actions and potential failures. While comparisons between GNN-based and non-GNN RL methods were conducted, detailed validation of specific GNN architectures and alternative models was lacking. Another method, SMAAC/AS, attempted to improve by incorporating goal topology without afterstate representation, but its poor performance underscored the effectiveness of afterstate representation. YZM, another baseline from the same L2RPN challenge, initially showed promise by limiting the considered actions but struggled to further improve due to the challenge of exploring a vast action space. SMAACsignificantly outperformed all other methods, highlighting its effectiveness in learning and managing complex power grids. Additionally, the study explored how different low-level control rules impact overall performance, finding that the high-level SMAAC policy can adapt to various low-level rules as long as they partly achieve goals.
In Sar et al., (2023), two model-free algorithms ( Soft Actor-Critic Discrete (SACD), PPO) are extensively evaluated and compared in their independent and dependent variants within the multi-agent framework. Independent versions have each agent with its own critic, actor, and replay buffer, but struggle with coordination and stability. Dependent versions use a centralized critic to improve coordination among agents, enhancing stability and performance. SACD achieves optimal scores in its single-agent form but is unstable in multi-agent setups, except for DSACD with tuned parameters. PPO, both single- and multi-agent, converges effectively, with single-agent PPO showing faster convergence. Moreover, PPO shows less sensitivity to hyperparameter changes in the multi-agent setting. The model architecture, particularly the GNN employed is not further validated against feedforward networks or other GNN architectures.
Zhao et al., (2022) employs GraphSAGE networks that are evaluated on a modified IEEE 118-bus system. The study includes training these networks unsupervised under specific scenarios and testing them on unseen scenarios with varying grid topologies. They use 2D t-SNE visualization to show the network’s robust representation of the power grid across different setups. Comparing their GraphSAGE-based DRL framework with a dense-based PPO algorithm, they find that GraphSAGE maintains high performance even with changing grid structures, unlike the dense-based approach, which struggles to adapt effectively. However, the evaluation primarily focuses on training outcomes rather than specific power grid performance metrics such as survival time of the agent.
In Xu et al., 2022a , their simulation-driven GRL method is evaluated using the L2RPN Robustness Track challenge dataset. The method shows improved performance over training episodes, combining decisions of sub-agents for collaborative decision-making. Compared to a "do-nothing" approach, it effectively prevents overloads across various scenarios. The model demonstrates rapid decision-making with an average time of 35 ms per time step. They analyze GAT and GCN-based models, finding that GAT supports stable power grid operation better, offering improved stability and economic benefits. However, a comparison with feedforward neural networks is missing. Their Long-Short-Term Action Deployment Strategy outperforms fully reward-guided and enumeration strategies by managing overloads effectively with fewer actions. The chosen action threshold of $0.98$ is validated to yield optimal results.
Wu et al., 2023b evaluated their GRL approach for managing BESS in against baseline methods like DQN and DDPG. They also compare different neural network architectures, focusing on the advantages of their spatio-temporal Chebyshev GCN (Cplx-STGCN) over feedforward, convolutional, and recurrent networks for extracting key features in power grid management. Additionally, the study tests two hybrid OPF solver methods, DeepOPF and DC3. They compare these methods with existing RL approaches like TD3, assessing metrics such as testing rewards and control over power generation and voltage magnitude. Overall, their constrained GRL framework shows superior performance compared to traditional optimization methods and existing RL techniques in managing BESS in power systems.
4.5 Discussion
The analyzed approaches maintain consistency in state representation aligned with the grid2op environment, yet they vary in the specific information included. For example, Xu et al., (2020) integrates extensive grid features like topology, generator characteristics, and power flow, while others such as Yoon et al., (2021) focus on afterstate representations to capture deterministic changes post-action. This adaptability of GRL methods reflects their ability to handle diverse inputs within the framework.
Generally, the presented approaches heavily rely on the grid2op framework Donnot, (2020) for training and testing, which is designed by transmission grid operators but doesn’t fully model all real-world grid aspects. Consequently, the rewards, actions, states, and graph representations are constrained by grid2op functionalities. Notably, while graph representation significantly affects GNN performance, there’s a lack of analysis on optimizing grid representations within these frameworks.
The diversity of RL methods is evident, with each approach employing tailored algorithms for complex power grid control. For instance, Xu et al., (2020) uses Double-Q-learning with soft constraint enforcement, while Taha et al., (2022) and Xu et al., 2022a employ MCTS for optimal action sequencing. To manage expansive action spaces, techniques like afterstate representations and expert knowledge-based action reduction are utilized to avoid unnecessary interventions in stable grid conditions. However, it is noteworthy that generally there is a limited application of model-based techniques, such as in Fuxjäger et al., (2023). Notably, no model-based technique employing GNNs has been developed so far. Similarly, it is important to highlight that state-of-the-art model-free RL methods, such as BBF Schwarzer et al., (2023), have not been applied.
Exploration poses challenges in environments with strict physical constraints, often leading to violations. Strategies like simulating multiple actions or using surrogate models for faster predictions mitigate these risks. Rewards vary across approaches, reflecting trade-offs in optimizing multiple objectives, with RL enabling comprehensive integration of network control aspects despite potential trade-offs between conflicting objectives.
GNNs are pivotal for extracting features from power grid topologies, enhancing convergence and generalizability across different grid configurations. While GraphSAGE, GATs, and transformer-based models are commonly employed, their evaluation against alternative architectures or feed-forward networks sometimes lacks depth, suggesting opportunities for further comparative studies to ensure robustness and applicability.
Challenges in evaluating GRL approaches include methodological diversity, the need for comprehensive validation, and scalability concerns. Some studies offer thorough evaluations of RL algorithms and architectures, while others focus more narrowly on effectiveness against specific baselines. Additionally, the widely used grid2op environment exhibits significant stochasticity that notably affects agent performance. Lehna et al., (2023) demonstrates performance variation across different random seeds and advocates for averaging results over multiple seeds (e.g., 30) to ensure reliability. However, some approaches discussed overlook this necessity, which could result in incomplete conclusions.
In terms of applicability, it should be noted that the sizes evaluated here are typically smaller than real-world grids, even for the larger test grids. Therefore, the proposed approaches are not yet scalable to real-world grid operations and serve more as a proof of concept. However, they pave the way for GRL-based decision support for grid operators.
5 Graph Reinforcement Learning for Distribution Grids
Distribution grids, which deliver electricity from the transmission grid to end users, are significantly impacted by the energy system’s transformation. Power generation has increasingly shifted from the transmission system to the distribution side Beinert et al., (2023) due to the rise of distributed renewable energy sources like photovoltaics. This shift causes voltage fluctuations that can threaten grid stability, as system voltages must remain within certain limits for safe operation. New consumers, such as distributed storages and electric vehicles, further exacerbate these fluctuations. Voltage control addresses these issues by flattening voltage profiles and reducing network losses using devices like voltage regulators, switchable capacitors, and controllable batteries Fan et al., (2022). Alternatively, topology control can reduce the need for additional online voltage control devices Xu et al., 2022b . Traditional voltage regulation devices are not designed for highly variable conditions and wear out quickly with frequent use. Therefore, distributed energy resources often include smart inverters that regulate voltages at a fast timescale.
These opportunities encourage the development of novel voltage control methods Srivastava et al., (2023). While DRL has shown promise in this area Duan et al., (2020), the use of GNNs is still emerging. We point out that DRL methods serve as a proof-of-concept and are still very far from practical deployment. Since voltage control involves sequential decision-making to determine optimal voltage set points, DRL is seems promising. Feed-forward neural networks struggle with the topological changes in distribution grids, making the combination of DRL and GNNs a potential better choice to capture the grid’s inherent graph structure.
Table 2: Overview of GRL approaches proposed for voltage-control applications in the distribution grid. The action column lists devices or variables modified by predicted actions (q for reactive power, p for active power). Grid sizes refer to the number of buses evaluated. The focus highlights key aspects or major differences from other approaches.
Model RL Action GNN Grid Size Focus/ Unique Feature
Operation control Yan et al., (2023) MAAC q(PV inverters) spectral GCN IEEE 141 zoned grids, primal- dual approach
Mu et al., (2023) MAAC q(PV inverters) spectral GCN IEEE 141, 39 zoned grids
Lee et al., 2022b PPO ESS, voltage regulators, capacitors GAT IEEE 13, 34, 123, 8500 graph augmentation, local readout func.
Wu et al., 2023a AC q (PV inverters) custom GSO (spectral) IEEE 33, 25 two-stage, hybrid (optimizer + RL), grid-specific filter in GNN
Wu et al., 2022c PPO q (PV inverters) custom GSO (spectral) IEEE 33,119 grid-specific filter in GNN
Wang et al., (2023) Multi-agent PPO ESS, p (generators) spatial GCN IEEE 69 , 123 real-time operation, multiple microgrids
Cao et al., (2023) AC q (PV inverters), ESS, var compressors GAT IEEE 33 surrogate GNN (grid embedding + reward)
Li et al., (2023) PPO p (generators) spatio-temporal GAT IEEE 33, 69, 118 consider temporal information
Xing et al., 2023a DDPG p (generators, PV, flexible loads), q (SVC) GAT IEEE 33, 119 consider many objective in reward
Xu et al., 2022b DQN grid topo spectral GCN IEEE 33, 69, 118 action space reduction via GNN + branch exchange
Emergency Situations Hossain et al., (2021) Double DQN binary load shedding spatial GCN IEEE 39 consider temporal information
Zhang et al., (2023) PPO binary load shedding, line switching graph capsule network IEEE 13, 34 stores graph neighbourhood statistics
Pei et al., (2023) double dueling DQN two-level load shedding GraphSAGE IEEE 39, 300 adaptability to unseen topologies
Zhao and Wang, (2022) Q-Learning reconnection of grid components GCN IEEE 123 and 8500 System restoration
5.1 Voltage Control
Voltage control is the key task in the distribution grid, managing reactive power set points to maintain grid stability. Reactive power supports voltage levels for delivering real power. The cost function typically includes indicators like power losses and congestion Srivastava et al., (2023). Keeping voltages within limits is crucial to avoid system damage.
The AC power flow in a grid is modeled by highly non-linear equations, making the optimization problem non-convex. To simplify, it is often linearized using methods like semi-definite programming or second-order case programming. For exact models, numerical methods like Newton-Raphson or Gauss-Seidel are used. Heuristics, such as particle swarm optimization, address the non-convexity of the problem formulation Srivastava et al., (2023). Deep learning approaches offer an effective alternative since they suited for non-linear problems, overcoming the limitations of traditional methods in handling the complexity and dynamics of smart grids.
In this work, we distinguish between two cases of voltage control: operation control and emergency mode. In operation control case, i.e. a stable grid state, voltage control is typically addressed via the control of reactive power. In contrast, emergency situations require a more drastic measures such as load shedding, i.e., the cutting of loads to prevent the grid from blacking out.
5.1.1 Operation Control
Recent studies show that the application of GRL in voltage control for distribution grids. These studies highlight the challenges posed by high levels of renewable energy integration, aiming to minimize network loss, mitigate voltage oscillations, and keep operational costs low. Reactive power control of generators and loads is widely used to manage bus voltage, directly influenced by local reactive power. With future inverter-based generation and digitized grids, maintaining voltage within the desired range (e.g., $0.9$ to $1.1$ p.u.) can be achieved through reactive power control.
RL Framework
Rewards
Most approaches in this area utilize a similar reward function to encode these objectives. Grid stability hinges on maintaining voltages within specified limits, leading these approaches to penalize voltage deviations from a set reference value. Wu et al., 2023a and Wang et al., (2023) focus solely on these voltage-based rewards. Many methods combine penalties for voltage deviation with other terms in a weighted sum to address multiple aspects supporting grid stability and cost minimization. For instance, Yan et al., (2023) and Mu et al., (2023) incorporate penalties for overall power loss, while Lee et al., 2022b add penalties to discourage frequent control changes that could cause equipment wear. Mu et al., (2023) include a voltage barrier function to constrain voltage ranges.
Meanwhile, Wu et al., 2022c and Li et al., (2023) integrate penalties for PV curtailment and voltage oscillation, with Li et al., (2023) also considering renewable energy integration and generation costs. Cao et al., (2023) base their reward on voltage deviation and use a surrogate model to estimate voltage and power loss. Balancing these reward components is crucial as optimizing one can affect others negatively, often requiring thorough parameter search.
Although Xu et al., 2022b focus on grid stability through changes in power grid topology, they too base rewards on voltage offsets and constraint violations. They introduce a penalty for recommending actions on already disconnected lines and define rewards based on tree metrics to ensure topological actions adhere to certain properties, such as avoiding loops or disconnected
Actions
Voltage regulation primarily involves adjusting actuator set points. Approaches proposed by Yan et al., (2023), Mu et al., (2023), Wu et al., 2022c , Cao et al., (2023), Wu et al., 2023a , and Xing et al., 2023a adjust reactive power outputs of PV inverters. Xing et al., 2023a also control the active power of Energy Storage Systems (ESS) and flexible loads, along with reactive power of static var compensation. Cao et al., (2023) additionally predicts set points for static var compressors. Some methods, like Lee et al., 2022b , Wang et al., (2023), and Cao et al., (2023), learn to manage ESS. Lee et al., 2022b also consider modifying configurations of voltage regulators and capacitors.
Alternatively, adjusting active power of generators, including renewables like wind and PV (Li et al., (2023), Wang et al., (2023)), or diesel generators (Wang et al., (2023)) alongside reactive power, is another option. For instance, Li et al., (2023) manipulate active power and voltage magnitude of various generators. In contrast, Xu et al., 2022b focus on modifying grid topology by disconnecting and reconnecting lines.
States
GRL approaches predict actions based on states such as voltage measurements, load demand, and power generation (both active and reactive). Methods like Lee et al., 2022b , Xu et al., 2022b , and Wu et al., 2023a also consider the status of actuators such as ESS or tap changers. In contrast, Wang et al., (2023) incorporates electricity grid prices. All approaches typically begin by embedding the distribution grid’s current state using a GNN, which encodes feature vectors of states and the grid’s topology represented by the adjacency matrix. These encoded representations are then used by the RL algorithm for decision-making.
RL Algorithms
The RL algorithms used in these studies primarily include Actor-Critic, PPO, and Q-Learning paradigms, each tailored to specific setups and objectives. For instance, Lee et al., 2022b , Wu et al., 2022c , Xing et al., 2023a , and Li et al., (2023) adopt a straightforward approach using GCNs to embed the distribution grid and train policies. Li et al., (2023) and Xing et al., 2023a employ DDPG for training, while Lee et al., 2022b and Wu et al., 2022c utilize PPO agents. Xing et al., 2023a further emphasize computational efficiency for real-time scheduling.
Apart from single-agent solutions, Yan et al., (2023) and Mu et al., (2023) address zoned distribution networks using a multi-agent Actor-Critic approach, where each agent manages a feeder. Yan et al., (2023) employs a centralized training approach with decentralized decision-making, ensuring physical constraints via a primal-dual method similar to that of Wu et al., 2023b , as described in Sec. 4.2. Similarly, Mu et al., (2023) integrates GCNs into the critic for modeling agent interactions. Wang et al., (2023) employs a multi-agent PPO setups for a distribution grid with multiple microgrids, where each is controlled by an agent that manages its power schedule.
On the other hand, Wu et al., 2023a proposes a two-stage approach: day-ahead optimization using Mixed Integer Second Order Cone Programming, followed by actor-critic learning for voltage regulation in a continuous action space with GCN-enhanced grid embeddings. Cao et al., (2023) also employs a GCN but follow it with a fully connected deep autoencoder for feature dimension reduction in an actor-critic framework, focusing on mapping grid representations to control actions.
In contrast, Xu et al., 2022b manages voltage control by adjusting grid topology. They address the NP-hard complexity, which grows exponentially with grid size, using a method to reduce the action space. They employ deep Q-Learning with a GCN-based Q-network to predict which line to disconnect. For reconnection, a branch exchange based on graph theory is used, adhering to radial constraints that ensure a tree-like network structure without loops. his approach involves selecting lines for reconnection from a basic cut set to maximize rewards, effectively reducing the action complexity.
Graph Embeddings
The GNN architectures in GRL vary significantly across the presented approaches. These GNNs primarily extract features from graphs representing power grids, where nodes correspond to buses and edges to power lines. Nodes include grid information such as voltage, active and reactive power, and the status of controllable actuators (voltage regulators, capacitors, batteries).
For instance, Lee et al., 2022b use a GCN in a straightforward approach with a PPO algorithm and augment their graph with additional edges to expedite information propagation. Similarly, Xing et al., 2023a employ a GAT (cf. Eq. 3) to capture topology branch correlations and power information and trained with DDPG.
Xu et al., 2022b utilize deep Q-Learning with a GCN-based Q-network to predict the best line for disconnection, using spectral convolutions (cf. Eq. 5) for feature extraction. Similarly, Yan et al., (2023) and Wu et al., 2022c apply spectral graph convolutions to merge topology and feature information in Actor-Critic setups. The spectral convolutions act as a low pass filter that suppresses noise in the input data, while the message passing aggregates the features of neighbouring nodes and fills in missing values. Wu et al., 2022c further include a graph shift operator based on AC power flow equations, embedding voltage angle and magnitude. This operator, combined with learnable weight matrices, improves grid feature extraction. They also incorporate temporal information by aggregating node embeddings from the past 10 time steps. Wu et al., 2023a and Li et al., (2023) also introduce novel graph shift operators in their GCNs, with Li et al., (2023) incorporating spatio-temporal attention mechanisms to capture both spatial influences and temporal dependencies across different resolutions.
On the other hand, Mu et al., (2023) employ a tree structure to represent the grid for their multi-agent actor-Critic approach, where a 2-layer spectral GCN integrates information from neighbouring agents. In another multi-agent study, Wang et al., (2023) use a GCN with an actor-critic approach to extract critical network topology and system information for decentralized microgrid voltage regulation. They select critical buses based on their connection to generators, microgrids, or feeder endpoints. These critical buses form the nodes in the grid’s graph. Electrical distance, calculated from the sensitivity coefficients derived from the Jacobian matrix, identifies connections between critical buses. This information is integrated into the adjacency matrix. The GCN, consisting of two graph convolutions, extracts features from the distribution grid, followed by max pooling and two feed-forward layers to learn a control policy.
As an alternative to directly learning the GNN weights using the RL loss, Cao et al., (2023) use a supervised GCN trained on historical power flow data to predict node voltages, which are then used in an actor-critic algorithm for feature extraction without updating the model weights.
Experiments and Evaluation
Most approaches evaluate their performance on IEEE grids with a size ranging from 5 to 300 buses. The authors evaluate the presented approaches with metrics like voltage deviation, network energy loss, or voltage violation rates. Comparisons typically include traditional optimization methods and other deep RL approaches as benchmarks.
The advantage of GNNs over dense-based RL agents is demonstrated in many of the approaches analysed. Starting with Lee et al., 2022b , the authors tested their method on PowerGym Fan et al., (2022) grids with 13, 34, 123, and 8500 nodes. Comparing the performances, their GNN-based PPO method slightly outperformed the dense PPO on smaller grids but converged slower on larger ones due to message-passing nature of the GNN layers. Additionally, experiments with missing or noisy observations showed that the superiority of the GNN-based approach in terms of robustness. Moreover, the paper finds voltage regulators impact the grid globally, while batteries and capacitors have local effects. To address this, they add edges between nodes with voltage regulators and use a local readout function for the controllable nodes, improving the robustness and performance of the GNN-based PPO approach. Similar to Lee et al., 2022b , Wang et al., (2023) compared their graph-PPO with a dense-PPO method on IEEE grids of 33 to 123 buses, demonstrating near-optimal performance and a better scalability.
Similar grid sizes are addressed by Li et al., (2023), who significantly outperform optimisation-based benchmarks. Their GRL approach achieves faster inference times, higher rewards, lower voltage fluctuation rates, and higher renewable accommodation. Furthermore, is has a faster convergence due to the spatio-temporal attention in their GNN architecture. Their analysis of the attention masks shows strong connections to high-power buses and that 10 time steps are sufficient to capture necessary temporal information, as correlation weakens beyond this.
Yan et al., (2023) compare their primal-dual GRL model with single-agent and multi-agent DDPG methods and their primal-dual variants. The GRL method effectively minimizes network energy losses and voltage deviations, even with noisy or incomplete data. Traditional volt-var optimization, while performing best with its access to global information, struggles with noisy or missing data. Similarly, Mu et al., (2023) find their GNN-based approach achieves more stable voltages and fewer violations compared to optimization methods and multi-agent DDPG on 33- and 141-bus grids. Further, it showed robustness to line and bus deletions. Wu et al., 2023a also show improved voltage profiles and reduced power losses on modified IEEE 33-node and 25-node systems compared to dense-based, CNN-based DDPG, and conventional optimization methods.
Wu et al., 2022c focused on different scenarios of outage attacks in 33- and 119-node systems with one-year smart meter readings and PV generation of a real distribution system. They showed that their approach mitigates oscillations from cyber-attacks even when 50% of inverters were corrupted, though they did not benchmark against other methods. Therefore, these results are difficult to interpret. Particularly, the utlilization of GNNs remains to be validated.
Xu et al., 2022b evaluated their results only against algorithmic solutions, showing their method was faster than the heuristics and close to optimal, with improvements from using GCN, branch exchange, and action separation. In an ablation study, they evaluate the benefits of the individual components of the proposed framework.
Finally, Cao et al., (2023) conducted comparative tests on the IEEE 33-node system, evaluating their physics-informed GAT-SAC approach against several control methods, including SAC variants and GCN-SAC. Their method outperformed others in reducing voltage deviations and maintaining safe voltage levels, especially under noisy conditions. Ablation studies emphasized the importance of the GAT-based network and the added robustness from the deep autoencoder. Tests on the IEEE 119-node system confirmed the method’s scalability and effectiveness in larger networks.
5.1.2 Distribution Grid Control in Emergency Situations
Another way to control grid voltages is load shedding, which involves deliberately disconnecting certain loads. This drastic measure is usually a last resort to avoid total blackouts. Conversely, in case of a full or partial blackout occurs, system restoration is needed to restart the grid. This involves coordinated steps to reconnect loads, restore generation capacity, and ensure the integrity of transmission and distribution networks. In the following, we present three load shedding and one system restoration approach that utilize GRL.
Load Shedding
The goal is to develop a GRL agent that learns to make load shedding decisions, as proposed by Hossain et al., (2021), Zhang et al., (2023), and Pei et al., (2023), which present three different GRL approaches for handling critical system loads.
The problem is modeled as an MDP, predicting actions based on grid state observations like power demand, generation, voltage measurements, and topology. Pei et al., (2023) further include historical node voltages. As for the decision-making, Hossain et al., (2021) and Zhang et al., (2023) model decisions as binary actions (load shedding or not), while Pei et al., (2023) considers shedding 5% or 10% of the load. All three approaches focus on heavy load nodes, with action space scaling accordingly. Zhang et al., (2023) also includes switching lines on and off. The reward in all approaches aims to maintain stable voltage levels and prevent system collapse. For instance, Hossain et al., (2021) gives a high negative reward if voltage stays below 95% of the nominal after 4 seconds, and a positive reward for voltages within predefined levels. The reward also minimizes load shedding or maximizes power supply (Zhang et al., (2023)) and penalizes actions violating system constraints.
RL algorithms used include DDQN with an $\epsilon$-greedy strategy and experience replay (Hossain et al., (2021) and Pei et al., (2023)), with Pei et al., (2023) also using a dueling architecture. Zhang et al., (2023) uses PPO with a hybrid policy network combining a fully connected network and a GCN.
In all papers, the distribution grid is modeled as a graph, with buses as nodes. This includes elements like substations, loads, and generators that connected by edges representing power lines and transformers. Node features include grid state data such as voltage measurements and loads. All approaches use a GCN for feature extraction, but with different architectures. Hossain et al., (2021) uses a simple GCN with three graph convolution layers, followed by a fully connected layer. Zhang et al., (2023) employs a graph capsule network (cf. Sec. 3) that embeds whole graphs and stores additional neighborhood statistics to avoid information loss. A node embedding then contains higher-order statistics for each feature within the local neighbourhood as well as compressed node labels, similar to Weisfeiler-Lehman colourings Verma and Zhang, (2018). Conversely, Pei et al., (2023) uses GraphSAGE, which samples and aggregates the local neighborhood of nodes.
The presented papers conduct experiments on IEEE case example grids. Hossain et al., (2021) and Pei et al., (2023) use the 39-bus system, while Zhang et al., (2023) tests on the modified IEEE 13-bus and 34-bus systems. Each method trains on different topological configurations with random fault locations.
All three approaches outperform feed-forward neural networks in convergence and average reward, especially on unseen topologies. Hossain et al., (2021) evaluates on 32 topologies. Pei et al., (2023) also tests their GNN-based method on the 300-bus system, managing larger action spaces better than feed-forward networks. Additionally, Zhang et al., (2023) and Pei et al., (2023) benchmark against traditional optimization techniques, achieving near-optimal performance with faster computation. Pei et al., (2023) demonstrates superior adaptability and efficiency, favoring GraphSAGE over a classical GCN. These methods enhance voltage control under varying grid topologies, showing the robustness and efficiency of GNNs compared to conventional methods.
System Restoration
On a different note, in case earlier mitigation actions fail and a blackout occurs, rapid system restoration is crucial to reconnect loads and restart the grid promptly. Zhao and Wang, (2022) approach this optimization with a routing model, treating generators as energy depots to connect to as many loads as possible. They employ a multi-agent approach where a Q-Network guides actions, using an encoder to observe generator capacities, switch statuses, and load conditions. A GCN extracts features from local and neighboring agents for the Q-Networks. This method outperforms single-agent DQN and other multi-agent baselines using feedforward networks, as well as CPLEX, a mathematical optimization method, in terms of accuracy and speed. Case studies validate this approach on IEEE 123 and 8500 node test systems Zhao and Wang, (2022).
5.2 Other Use Cases
Two additional use cases for distribution grids include loss minimization and economic dispatch. Loss minimization focuses on optimizing the power grid state by adjusting topology or generator set points to reduce losses in branch elements. Economic dispatch, on the other hand, aims to minimize the operational costs of generators.
Jacob et al., (2022) focuses on Distribution Network Reconfiguration (DNR) to minimize losses and enhance resilience by dynamically reconfiguring the network’s topology through sectionalizing and tie line switches. This adapts the distribution network to changing conditions like demand fluctuations or disruptions. Their states include the current network topology, status of connections (open or closed), power demands at buses, generation outputs, bus voltages, branch currents, and total network loss for efficiency assessment. Rewards in Jacob et al., (2022) incentivize loss reduction and compliance with topological constraints, penalizing actions that disconnect loads or alter the radial structure, while rewarding feasible network exploration
Meanwhile, Chen et al., (2023) optimizes economic dispatch in power systems with high renewable energy penetration for economic efficiency and stability under dynamic and uncertain conditions, involving adjustments in generation, renewable output, and ESS charging/discharging power. The states include load demands, generation and RES outputs, ESS state of charge, and the current timestep, offering a comprehensive view of the power grid’s dynamic state. Rewards are designed to optimize economic dispatch and system stability, combining economic operation costs with a stability component addressing voltage violations, encouraging actions that balance efficiency and stability.
Regarding the graph representation, the nodes in Jacob et al., (2022) represent substations and buses connected by edges for power lines and transformers. They employ a Capsule-based Graph Convolutional Network (GCAPCN) (cf. Sec. 3) to capture both local and global graph features. Likewise, Chen et al., (2023) adopts a power system graph representation where nodes represent buses and edges represent power system branches. They utilize a GCN to process node features including load demand, generation outputs, RES outputs, ESS state, and timestep information, mapping the system state to an action policy through layers of transformation and feature extraction.
Table 3: Overview of other relevant GRL approaches proposed for distribution grids. The RL algorithm, the action as well as the GNN type are specified. The focus defines the objective of the proposed approaches.
Model RL Action GNN Focus
Other approaches Jacob et al., (2022) PPO Line switching GCAPCN Loss minimization with DNR
Chen et al., (2023) SAC Adjusting generation, RES output, ESS charging/discharging power GCN Economic dispatch
For loss minimization, Jacob et al., (2022) use the PPO algorithm for training. The GCAPCN encodes grid details such as losses, bus voltages, branch currents, and topology for the policy network, while a feedforward neural network serves the value network. In contrast, Chen et al., (2023) apply the SAC algorithm with the GCN layers in both the actor and critic roles to optimize dispatch policies off-policy.
The approaches are evaluated as follows: Jacob et al., (2022) compare their GCAPS-RL method against two conventional baseline methods, Mixed-integer Second-order Conic Programming (MISOCP) and binary particle swarm optimization (BPSO), as well as a feed-forward approach to assess the benefits of GNNs. Experiments are conducted on modified IEEE distribution test networks (13-bus and 34-bus systems) using power flow simulations. GCAPS-RL demonstrates superior real-time decision-making and adherence to topological constraints, outperforming the feed-forward counterpart particularly in the 34-bus system.
Meanwhile, Chen et al., (2023) perform case studies on a modified IEEE39 power system configuration with conventional generators, renewable sources, and an ESS. Their GRL approach is compared against Optimal Solution with Perfect Information (OSPI) and Model Predictive Control (MPC) policies with varying prediction horizons, as well as feed-forward SAC. GRL exhibits strong convergence, effective policy performance, and superior scalability when the system topology changes. It achieves near-optimal results and reduces costs compared to all baselines, with efficient computation meeting practical requirements, especially under extended prediction horizons.
5.3 Discussion
Voltage control and grid control in emergency situations using DRL techniques encompass a variety of approaches and considerations. One key aspect is the design of reward functions, which typically address voltage deviations but may also incorporate additional factors such as the integration of renewables or power loss. However, balancing these objectives poses a non-trivial challenge as already mentioned in Sec. 4.5.
In terms of control devices, most methods focus on PV inverters or other actuators, though some also consider ESS or adjust the active power of generators. Notably, Xu et al., 2022b stands out for its consideration of topological actions, offering a unique perspective in this domain.
The control strategies for distribution grids that are divided into zones or microgrids involve multi-agent setups, where each agent corresponds to a distinct zone. Here, the treatment of global system knowledge varies, with approaches either training globally and evaluating locally or utilizing a centralized critic for action evaluation. In this setup GNNs can be used to fuse the information from different agents as presented in Mu et al., (2023).
The implementation of GNNs varies widely across studies, with no consensus on whether to implement them in the actor, critic, or both. To map node embeddings to actionable vectors, facilitating dimensionality reduction and action prediction, the readout methods vary, ranging from fully connected neural networks to autoencoders to 1D convolutions. Weight sharing between surrogate models and actors or critics is observed in some approaches. While graph representations remain largely consistent, GNN architectures exhibit significant diversity. Spectral GCNs are prevalent, although less common in other domains. This might be due to the fact that spectral convolutions come from signal processing and thus are closer to electrical engineering. Some approaches also utilize temporal information by including historical data, but most approaches consider static information only. Multiple studies introduce customized Graph Structure Operatorss which make use of domain knowledge about certain properties of the system. The respective approaches show that the feature extraction benefits from the incorporation of known system dynamics. Domain knowledge can also be used for the augmentation of the graph representation of grids to define edge weights and determine critical connections in beforehand.
Also already pointed out in Sec. 4.5, GNNs have outperformed fully connected neural networks in terms of transferability and adaptability to topology changes. Experiments with deleted grid elements or completely different grid structures can be handled without a drastic drop in performance, which is not the case for fully connected neural networks.
The ability of GNNs to handle noisy or missing data effectively is frequently cited as a motivation for their use. In the according experiments including generator failures, deleted lines or nodes, all works demonstrate the robustness of GNNs. In real grid operation, faulty sensor data or unknown parameters are the normal case rather than an exception, therefore this is a clear advantage.
The experiments and evaluations of these approaches encompass a wide range of considerations. Most studies conduct experiments on IEEE grids, with grid sizes varying from small-scale systems with as few as 5 buses to large-scale networks with up to 8500 nodes. The grid sizes exhibit considerable variation, impacting the scalability and generalization of approaches. Notably, some methods demonstrate effective performance even without access to global information, underscoring the robustness of certain DRL strategies in this context. Evaluation metrics typically include voltage deviation, network energy loss, or voltage violation rates, reflecting the overarching objective of voltage control and load shedding. Traditional optimization techniques and heuristics serve as benchmarks in many studies, highlighting the comparative performance and efficiency gains of DRL-based approaches. Additionally, comparisons with other DRL methods, including dense-based and CNN-based approaches, shed light on the relative advantages of graph-based methodologies. The experiments often demonstrate the scalability, stability, and real-time deployment capabilities of the proposed frameworks, showcasing their effectiveness in addressing complex power system control challenges.
6 Other Applications
Table 4: Overview of other relevant GRL approaches.
Model RL Action GNN Focus
Energy Market Rokhforoz et al., (2023) Multi-agent Actor-Critic Pricing GCN traditional electricity market
Lee et al., 2022a DQN, DRQN, Bi-DRQG, PPO buy/sell/hold GCN, Bi-LSTM P2P power trading, maximizing integration of renewables
Communication
Islam et al., (2023) Q-Learning Routing, set queue service rate spectral GCN reduce end-to-end latency, packet loss rate
EV charging Xu et al., 2022c Double prioritized DQN($\lambda$) CS recommendation GAT Combine RL and Dijkstra for CS recomm. and rounting
Xing et al., 2023b Rainbow DRL CS recommendation, selection of road segments GAT Bi-Level RL for CS recomm. and routing
6.1 Energy Market
GRL opens new possibilities in the energy market, especially in decentralized bidding or direct trading among entities. Traditionally, bidding strategies are centrally managed, requiring full information on all generation units. This is often unavailable due to privacy concerns and results in large-scale, computationally expensive problems. Distributed decision-making, using multi-agent RL and GNNs has the potential to provide efficient and scalable solutions.
We limit our focus to the context of power grids and review two papers using GRL to optimize energy trading strategies considering grid topology. Rokhforoz et al., (2023) focuses on the traditional market where generation units set their prices, and a market operator optimizes bids for the lowest overall cost. In contrast, Lee et al., 2022a explores P2P trading, where individuals trade electricity directly, promoting renewable integration.
Rokhforoz et al., (2023) consider a two-level optimization: first, each unit sets a bidding price; second, the market operator determines the market price. The goal is to maximize each unit’s profit. They propose a multi-agent actor-critic algorithm, where each unit’s actor network selects a bidding strategy based on previous market data. After the market price is set, rewards are calculated. The critic network, a GNN, updates the bidding strategy based on the grid topology. Experiments on the IEEE-30-bus and 39-bus systems show that the GNN approach outperforms the baseline using an MLP-based critic, particularly with varying generation capacities. The GNN also demonstrated better transfer capability when tested across different systems.
In the approach by Lee et al., 2022a , energy is traded directly between prosumers without an intermediary market. The setup includes multiple nanogrids, an information network, and a business network for trading. The proposed RL algorithm learns trading strategies to minimize maximum load and maximize renewable integration, including power from discharging EVs. The agent’s actor and critic are hybrid models combining a GCN with a Bi-LSTM to process time series data on prosumer consumption and production. The model inputs include cluster demand, renewable supply, system price, and demand response. Actions are buy, sell, or hold, and rewards are based on a rule-based baseline. Multi-objective optimization includes load shifting. The authors compare various RL methods, including DQN, Bi-LSTM, and PPO, using a nanogrid with real usage data. The PPO GCN-Bi-LSTM approach achieves the lowest electricity cost and performs better than other methods, significantly reducing average electricity costs with P2P trading.
Discussion
Both studies demonstrate that GNNs are promising for optimizing energy markets, particularly as decentralized approaches gain popularity for computational and privacy reasons. GNNs allow consideration of neighboring market participants’ information without creating large-scale problems, unlike traditional deep learning methods that treat participants as independent samples. Experiments show that incorporating information from nearby nodes enhances overall market profit. GNNs improve both Actor-Critic and Q-Learning RL methods by capturing interdependencies missed by MLP-based methods, learning more representative grid embeddings crucial for RL decision-making. These findings highlight the potential of GRL in energy markets, with more GRL-based approaches expected in the future.
6.2 Power Communication Networks
Apart from the power transmission itself, modern energy systems also transmit information for monitoring and control, requiring efficient routing in communication networks to avoid critical information loss. Unlike physical power transmission (cf. Sec. 4 and Sec. 5), these networks operate on the cyber layer.
Islam et al., (2023) develop a prioritized routing strategy considering grid states. They address two packet types: periodic packets with fixed schedules and emergency packets needing low latency. The goal is to reduce end-to-end latency, packet loss, and unmet latency requirements using software-defined networks that adapt dynamically to grid conditions.
Two Q-Learning RL algorithms are trained: one for routing paths and another for queue service rates to minimize congestion. The first agent selects feasible paths, while the second predicts queue service rates at switches. Rewards are based on the difference between switch capacity and queue state in order to accelerate queue emptying. A GNN predicts future grid states to inform the queue service rate agent, though the GNN is trained separately from the RL agent. The model, using spectral GCN with Chebyshev polynomials, is trained on IEEE grid traffic data. Experiments on the cyber layers of the IEEE-14 and 39-bus systems show the approach’s effectiveness in managing grid communication through message exchanges between devices and control centers.
6.3 Electric Vehicles Applications
The rapid growth of electromobility has posed challenges with increasing demands for electrical energy and straining electricity networks, particularly with variable load behaviors and decentralized renewable energy supplies. Efficient infrastructure use has become critical. DRL has found applications across several domains including proof-of-concepts for charging management Bayani et al., (2022); Sadeghianpourhamami et al., (2020); Li et al., (2020); Silva et al., (2020), station recommendation Xing et al., 2023b ; Xu et al., 2022c , navigation Xing et al., 2023c ; Xu et al., 2022c ; Xing et al., 2023b , and pricing optimization Zhang et al., (2022). These applications optimize energy distribution, route EVs efficiently, and dynamically adjust pricing to balance operator revenue with infrastructure efficiency.
We focus on GRL approaches and highlight two studies that include power grid constraints in optimizing charging processes, while noting the broad applications of GRL in various EV charging scenarios.
Xu et al., 2022c address the demands on fast charging stations caused by the increasing number of EVs, which affect coupled power-transportation networks. They propose a multi-objective system-level method to dynamically allocate EVs to suitable stations using DRL. This method considers the interests of EV owners, charging stations, traffic networks, and power grids, ensuring real-time responsiveness and efficient allocation of fast charging resources. EVs send charging requests when their state of charge drops below a threshold, and an agent recommends a charging station and guides them using Dijkstra’s algorithm. The multi-objective function optimizes EV wait times, charging station service balance, traffic congestion, and power grid voltage deviation.
Similarly, Xing et al., 2023b introduce a Bi-Level GRL approach for charging and routing in Transportation Electrification Coupled Systems. They use a bi-level problem formulation, with both levels employing a GNN-based feature extraction layer and a Rainbow-architecture DRL block for decision-making. The upper level recommends a charging station, while the lower level handles EV travel and route selection. Unlike Xu et al., 2022c , this approach uses a DRL agent for route planning instead of a conventional algorithm.
Graph Embeddings
GNNs leverage the inherent graph structures of transportation systems and power grids. Xu et al., 2022c designs a graph structure that captures connections between charging stations, roads, and power nodes. Charging stations bridge the traffic network and power grid, connecting to traffic nodes (midpoints of road sections) and power nodes based on geographical and power supply relations. The adjacency matrix of the charging network represents connections between traffic nodes (TN), power nodes (PG), charging stations (CS), and EV nodes. A unified expression method with type-specific transformation matrices projects features into a shared space, and GATs extract meaningful features from neighboring nodes using attention mechanisms. These learned representations are integrated with EV features for input to a DRL agent.
Similarly, Xing et al., 2023b utilizies GATs and introduces an instantaneous adjacency matrix for connections among EVs, CSs, TNs, and PGs, with smaller matrices representing different relationships (e.g., EV-TN, CS-TN). A node feature matrix stores energy and information features, with different rows representing system elements.
RL algorithms
When recommending a fast charging station, the decision is quick, but actual completion occurs after the EV finishes charging, causing significant delays. This delay, coupled with system complexity and unpredictability, makes forecasting the entire charging process challenging, impacting standard RL methods. To address this, Xu et al., 2022c introduce the double-prioritized DQN($\lambda$) method, integrating $\lambda$-return and experience replay to handle action delay and uncertainty. They use a small cache in replay memory for efficient training, prioritize crucial transitions, and employ an attention-based method to construct a cache that maintains high-quality training transitions. Additionally, a strategy regulates boundary actions to prevent detrimental decisions.
In contrast, Xing et al., 2023b employ a bi-level approach to address the credit assignment problem. The upper level selects charging stations less frequently, focusing on choosing a target CS, while the lower level deals with path navigation by selecting road segments. Rewards in each level consider charging costs, battery loss, time allocation, energy consumption, travel time, and penalties for exceeding voltage limits. The upper level interacts with charging stations and power grids, while the lower level interacts with traffic nodes after determining the target station.
Experiments and Evaluation
Xu et al., 2022c use a joint power-transportation simulation platform with an IEEE 33-node distribution network and a 25-intersection traffic network to validate their method. They optimize traffic conditions, user experience, and power grid stability, demonstrating superiority over distance-based methods across scenarios, even under charging station queue limitations. Training a user-oriented graph DQN($\lambda$) agent shows significant long-term benefits and user experience improvement. Combining GATs and DQN($\lambda$) outperforms other models in training, confirming its effectiveness in complex decision-making. However, a comparison of an MLP-DQN($\lambda$) remains to be performed in order to determine the effect of the GAT representation without any effects from the long term $\lambda$-return.
Xing et al., 2023b evaluate their Bi-Level GRL approach for Electric Vehicle Fleet Charging Guidance using real-size transportation-electrification data. They achieve 10.08% cost reduction and 16.45% time savings for owners compared to realistic scenarios without guidance. Compared to traditional methods, their approach lowers average total cost by 8.96% (SDP) and 4.73% (DRL), highlighting its efficiency in reducing charging costs and travel time. They suggest learning the GNNs weights directly with RL for enhanced robustness and scalability.
Discussion
Xu et al. Xu et al., 2022c demonstrate GRL’s ability in dynamically allocating resources for fast charging stations, emphasizing real-time responsiveness and multi-stakeholder considerations. Their approach highlights sequential decision-making’s role in balancing objectives across transportation and power networks. In contrast, Xing et al. Xing et al., 2023b present a Bi-Level GRL approach, focusing on efficient charging and routing coordination through GNNs. Their method excels in reducing charging costs and travel time, demonstrating GRL’s potential in real-world scenarios. While both employ GAT architecture, further comparisons of GNN architectures and robustness to noisy data are essential for future research, in order to further GRL’s applicability in sustainable transportation and energy infrastructure optimization.
7 Conclusion and Outlook
The exploration of DRL and GRL approaches in addressing the challenges of power grid control, voltage management, and other related use cases such as EV charging management reveals significant advancements and promising avenues for future research and applications. Across the transmission grids, distribution grids, and EV charging domains, these approaches demonstrate innovative solutions that leverage machine learning techniques to optimize system performance, enhance resilience, and facilitate the integration of renewable energy sources and electric transportation.
The pivotal role GNNs in RL approaches cannot be overstated. GNNs serve as the backbone for feature extraction from complex network structures, enabling the effective representation of power grid topologies. Their ability to capture spatial dependencies and relational information among grid components facilitates informed decision-making in multiple grid control applications. Moreover, the versatility of GNN architectures allows for customization to specific application requirements, ranging from spectral and spatial GCNs, to GATs, to GraphSAGE and beyond. Despite their widespread adoption in the presented papers, further research is needed to systematically evaluate and compare the suitability of GNN architectures for various control tasks within the power system domain. GNNs have shown to achieve better results than fully connected neural networks on unseen grid topologies and differing networks sizes.
In real-world power systems, data quality issues such as noisy sensor measurements and missing data are common due to factors like sensor malfunctions, communication failures, and data transmission errors. These data imperfections can significantly impact the performance and reliability of RL models. GNNs excel at capturing complex spatial dependencies and relational information from graph-structured data, allowing them to effectively distinguish meaningful patterns even in the presence of noisy data. By leveraging the inherent connectivity and topology of the underlying network, GNNs can smooth out noise and extract robust representations that encode relevant features for decision-making tasks. While a few of the presented approaches validated their GRL methods against noisy measurements, more research is needed to establish robustness across diverse datasets and real-world conditions.
Although different GNN architecture have been used across the approaches, the models are mostly very shallow with only 2-3 layers. As pointed out in Ringsquandl et al., 2021b , the graphs structure of GNNs differs from standard benchmark datasets, which makes well-established models not fully suitable to learning tasks on power grids. In fact, they find out that much deeper GNNs perform better, since they can capture long-range dependencies in graphs with a low cluster and high diameter as it it the case for power grids. Therefore, deeper architectures are an interesting topic of research especially in case of large-scale grids.
As for the RL methods, the research landscape for power grid control reveals a lack of application of state-of-the-art RL innovations. On the one hand, despite few efforts utilizing MCTS, the application of model-based RL methods remains limited, further exemplified by the absence of GNN based approaches. On the other hand, contemporary model-free RL innovations such as BBF Schwarzer et al., (2023) have yet to be explored in this context, which underscores a significant gap in current research.
Another promising direction for future research is the incorporation of domain knowledge. In the discussed approaches, it had been used to identify relevant grid connections or components to reduce the actions space or augment the graph to introduce meaningful edge weights and enable faster information diffusion. Further, the customized graph shift operators proposed by several approaches have shown to improve the performance in GNNs. Making use of the known relations and coherence in power grids, therefore, should be investigated further. Additionally, further research is needed to explore the incorporation of physical properties, such as power flow constraints, into agent training. This could be achieved, for example, through the utilization of physics-informed GNNs.
Variation in reward functions reflect the trade-offs involved in optimizing multiple objectives and highlights the ability of RL to handle complex, multi-dimensional optimization problems. Multi-objective RL offers flexibility by allowing the incorporation of various objectives into the reward function. However, balancing conflicting objectives remains a challenge, as optimizing one objective may have adverse effects on others. In future research, there is a pressing need to explore approaches where agents are capable of providing multiple suggestions instead of just a single solution. This could involve presenting a range of solutions along the Pareto front, showcasing the trade-offs between conflicting objectives, or offering alternatives that prioritize different objectives. By offering multiple suggestions, agents can enhance the trustworthiness and transparency of their recommendations, allowing users to gain a deeper understanding of the decision-making process. Moreover, presenting a variety of options empowers users to make more informed decisions that align with their preferences and operational constraints.
In the transmission grid domain, many approaches focus on smaller grid sizes. There is potential to extend these techniques to larger grids. Investigating how these methods scale with grid size and complexity would be valuable for real-world deployment. Furthermore, the grid sizes evaluated in many of the presented papers are typically smaller than real-world grids, even for the larger test grids. Action space reduction, and exploration in constrained environments underscore the need for further methodological advancements, particularly as the power grid size increases. Handcrafted rules and rule-based algorithms have been successfully integrated into the different GRL solutions. The synergy between rule-based decisions and GRL methodologies showcases the importance of leveraging domain knowledge to achieve effective grid control.
Notably, it’s important to point out the inherent bias towards the grid2op framework, which provides the data, training and testing environments for many approaches. While grid2op incorporates many real-world aspects, there’s a need to optimize graph representations for grid processing, which can significantly influence the performance of GNNs. Addressing the stochasticity in the grid2op environment is also crucial, as it significantly impacts agent performance. Averaging over multiple seeds is essential for reliability, yet often overlooked in some approaches. Transitioning GRL approaches from simulation environments like grid2op to real-world grid operation poses significant challenges, including computational efficiency, scalability, and integration with existing control systems.
For the distribution grid, the discussion surrounding voltage control and grid management GRL techniques underscores the diverse approaches and considerations within this domain. A central challenge lies in the design of reward functions, which typically prioritize voltage deviations while also accommodating factors like renewable energy integration and power loss. The choice of actuators, primarily focusing on PV inverters but also considering Energy Storage Systems and generator adjustments, highlights the need for adaptable control strategies. Multi-agent setups for zoned distribution grids present an interesting avenue, with varying approaches to incorporating global system knowledge. The implementation of GNNs varies widely, reflecting the ongoing exploration of their effectiveness in handling noisy or missing data. Despite consistent graph representations, there is significant diversity in GNN architectures. Evaluations span a range of grid sizes and metrics, demonstrating the scalability, stability, and real-time deployment capabilities of GRL-based frameworks. Additionally, further investigation into the augmentation of graph structures and the integration of temporal information could yield valuable insights for improving the performance and robustness of GRL-based approaches in distribution grid control.
The transition to fully autonomous systems must be approached cautiously, with a focus on developing decision support systems that augment human expertise rather than replacing it entirely. Therefore, exploring methods for enhancing the explainability of DRL approaches to ensure transparency and trustworthiness in the decision-making process remains crucial for the future. Besides explaining model decisions, assessing model uncertainty, and verifying model behavior under various conditions can help enhance transparency and confidence in DRL-based grid operation. The integration of GNNs into RL algorithms holds significant promise, since leveraging the inherent graph structure of power systems topologies can offer a more intuitive understanding of complex interactions. Architectures like GATs introduce interpretability through mechanisms such as attention coefficients, enabling the quantification of the importance of neighboring nodes. This does not only help to enhance model transparency but also facilitates more informed decision-making within RL frameworks.
8 Acknowledgements
This work is supported by
  1. 1.
    Graph Neural Networks for Grid Control (GNN4GC) founded by the Federal Ministry for Economic Affairs and Climate Action Germany under the funding code 020E-100626677.
  2. 2.
    AI4REALNET has received funding from European Union’s Horizon Europe Research and Innovation program under the Grant Agreement No 101119527. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
  3. 3.
    Reinforcement Learning for Cognitive Energy Systems (RL4CES) from the Intelligent Embedded Systems of the University Kassel and Fraunhofer IEE.
GNN
Graph Neural Network
GCN
Graph Convolutional Network
GAT
Graph Attention Network
EV
Electric Vehicle
RL
Reinforcement Learning
DRL
Deep Reinforcement Learning
GRL
Graph Reinforcement Learning
XAI
Explainable AI
TSO
Transmission System Operator
DSO
Distribution System Operator
PPO
Proximal Policy Optimization
CPES
cyber-physical energy system
OPF
Optimal Power Flow
DQN
Deep Q-Network
DDQN
Double Deep Q-Network
CNN
Convolutional Neural Networks
A3C
Asynchronous Advantage Actor-Critic
DDPG
Deep Deterministic Policy Gradient
DPG
Deterministic Policy Gradient
PPO
Proximal Policy Optimization
SAC
Soft Actor-Critic
TRPO
Trust Region Policy Optimization
MCTS
Monte Carlo Tree Search
SACD
Soft Actor-Critic Discrete
STGCN
Spatio-Temporal Graph Convolutional Network
BESS
Battery Energy Storage Systems
SMAAC
Semi-Markov Afterstate Actor-Critic
ESS
Energy Storage Systems
GSO
Graph Structure Operators
MISOCP
Mixed-integer Second-order Conic Programming
BPSO
binary particle swarm optimization
GCAPCN
Capsule-based Graph Convolutional Network
OSPI
Optimal Solution with Perfect Information
MPC
Model Predictive Control
RES
Renewable Energy Sources
DNR
Distribution Network Reconfiguration
MARL
Muli-Agent Reinforcement Learning
MDP
Markov Decision Process
BBF
Bigger, Better, Faster